### v9

# VFEFL: Privacy-preserving federated learning against malicious clients via verifiable functional encryption

Abstract
Federated learning is a promising distributed learning paradigm that enables collaborative model training without exposing local client data, thereby protecting data privacy. However, it also brings new threats and challenges. The advancement of model inversion attacks has rendered the plaintext transmission of local models insecure, while the distributed nature of federated learning makes it particularly vulnerable to attacks raised by malicious clients. To protect data privacy and prevent malicious client attacks, this paper proposes a privacy-preserving Federated Learning framework based on Verifiable Functional Encryption (VFEFL), without a non-colluding dual-server assumption or additional trusted third-party. Specifically, we propose a novel Cross-Ciphertext Decentralized Verifiable Functional Encryption (CC-DVFE) scheme that enables the verification of specific relationships over multi-dimensional ciphertexts. This scheme is formally treated, in terms of definition, security model and security proof. Furthermore, based on the proposed CC-DVFE scheme, we design a privacy-preserving federated learning framework that incorporates a novel robust aggregation rule to detect malicious clients, enabling the effective training of high-accuracy models under adversarial settings. Finally, we provide the formal analysis and empirical evaluation of VFEFL. The results demonstrate that our approach achieves the desired privacy protection, robustness, verifiability and fidelity, while eliminating the reliance on non-colluding dual-server assumption or trusted third parties required by most existing methods.
I Introduction
Federated learning[23] is an advanced and promising distributed learning paradigm that enables multiple clients to collaboratively train models without sharing local data. It embodies the principles of focused data collection and minimization, thereby mitigating many of the systemic privacy risks and costs inherent in traditional, centralized machine learning approaches. With the enactment of privacy protection and data security regulations in various countries, as well as the increase in public privacy awareness, the application of federated learning has gradually expanded e. g., healthcare, finance, and edge computing.
Although the federated learning framework offers a degree of privacy protection for client data during training, it brings new privacy and security issues [20]. One primary concern is the potential for privacy leakage through the local models uploaded by clients. With the advent and ongoing advancement of attack techniques, such as model inversion attacks[33, 17, 21], submitting local models in plaintext form is no longer regarded as a secure method. Accordingly, obfuscation and encryption mechanisms have been adopted to address this issue. Another threat is attacks from malicious clients. Due to its distributed architecture, federated learning is susceptible to adversarial manipulation by malicious clients. A malicious client may make the accuracy of the global model drop rapidly by uploading malicious models, or destroy the decryption by uploading a malicious decryption key or a wrong ciphertext. The above issues are important for federated learning systems to achieve their core goal (training high quality models while preserving data privacy). Privacy-preserving mechanisms (such as encryption schemes) can potentially lead to difficulties in detection and verification, while detecting and verifying encrypted models of the clients may compromise specific information. Existing solutions [13, 27] cannot be deployed in the basic federated learning architecture (one server and multiple clients), and need help from two non-colluding servers or additional trusted third parties, which introduces additional limitations and consumption.
Therefore it is very interesting and challenge to design a federated learning scheme that realizes privacy protection and resists malicious client attacks in the basic federated learning architecture. This paper proposes a privacy-preserving federated learning scheme against malicious clients based on decentralized verifiable functional encryption. Our contributions can be summarized as follows:
  1. 1.
    We employ a verifiable functional encryption scheme to encrypt local models in the federated learning, ensuring data privacy and correctness during encryption and decryption. The scheme supports verifiable evaluation of relations over multidimensional ciphertexts, making it well suited for federated learning scenarios. Furthermore, we formalize its definition and security model, and reduce its security in the proposed model to well-established computational hardness assumptions.
  2. 2.
    We propose a privacy-preserving federated learning scheme named VFEFL that protects local model privacy and defends against malicious clients. Compared with existing approaches, it enables verifiable client-side aggregated weights and can be integrated into standard federated learning architectures to enhance trust. Theoretical analysis demonstrates that VFEFL provides strong guarantees in both privacy and robustness.
  3. 3.
    We implement a prototype of VFEFL and conduct comprehensive experiments to evaluate its fidelity, robustness, and efficiency. Robustness is evaluated under both targeted and untargeted poisoning attacks. Experimental results demonstrate that VFEFL effectively defends against such attacks while preserving model privacy.
II RELATED WORK
II-A Functional Encryption
Functional encryption (FE) [5, 6] is a paradigm supports selective computation on encrypted data, while traditional public key encryption only implements all-or-nothing decryption method. An authority in a FE scheme can generate restricted decryption keys to allow receiver to learn specific functions of the encrypted messages and nothing else. Goldwasser[18] first introduced multi-input functional encryption (MIFE), which is a generalization of functional encryption to the setting of multi-input functions. Michel Abdalla [1] constructed the first MIFE for a non-trivial functionality with polynomial security loss under falsifiable assumptions for a super-constant number of slots.
However, whereas the input can be large, the original definition of FE limits the source of the encrypted data to a single client, which is not compatible with a wide range of real-world scenarios where data needs to be collected from multiple sources. Multi-client functional encryption (MCFE)[14] was proposed to addresses this issue. Further, decentralized multi-client functional encryption (DMCFE)[14] eliminated reliance on a trusted party. In the DMCFE scheme instantiated over an asymmetric pairing group, each client independently samples $s_{i}\leftarrow\mathbb{Z}^{2}_{p}$, and all clients jointly generate random matrices $T_{i}\leftarrow\mathbb{Z}_{p}^{2\times 2}$ such that $\sum_{i\in[n]}T_{i}=0$. $s_{i}$ and $T_{i}$ forms the secret key held by each client. This interactive procedure allows clients to locally construct their own keys without relying on a trusted third party.
The decentralized architecture of DMCFE brings forth additional challenges posed by potentially dishonest clients, prompting the proposal of verifiable functional encryption (VFE) [2] as a solution. Nguyen et al.[25] proposed a verifiable decentralized multi-client functional encryption scheme for inner products (VDMCFE), along with its formal definitions and security proofs. This work extended the DMCFE scheme by constructing the Class Group-Based One-Time Decentralized Sum (ODSUM) and Label-Supporting DSUM (LDSUM) schemes, leveraging the Combine-then-Descend technique to improve the efficiency of correctness proofs for functional decryption key shares. Additionally, it incorporated zero-knowledge proofs such as range proofs to ensure the correctness of ciphertexts and the validity of the encrypted content.
II-B Federated Learning
TABLE I: Comparison with other schemes
Proposed Works Techniques Robustness Self-Contained Verifiability Fidelity
Biscotti[28] DP+$S$
SecProbe[32] DP
ESFL[24] DP
SecureFL[19] PLHE+$S$
AegisFL[13] HE
RVPFL[22] HE
FheFL[27] FHE
FEFL[12] 2DMCFE
BSR-FL[31] NIFE
Ours $C$-DVFE
Note: Self-contained means the scheme works in a single-server setup without trusted third parties. Verifiability ensures the client follows protocol specifications. Fidelity [8] means no accuracy loss when no attack exists.
Federated Learning against Malicious Clients. Robust aggregation rules are one of the most common means of resisting malicious clients in federated learning. Next, we review some of the existing Byzantine robust aggregation rules (ARG).
  • FedAvg[23]. FedAvg is the most basic aggregation rule constructed when federated learning was first proposed. It directly computes the weighted sum of all local model to obtain the global model, where the weight of each client is determined by the proportion of its local dataset size to the total dataset size. However, FedAvg is not resistant to malicious client attacks, where the adversary manipulates the global model at will through submitted local models.
  • Krum[4]. This AGR selects global model update based on the squared distance score. Formally, for client $i$, the score is defined as $s(i)=\sum_{i\to j,n-f-2}\|W_{i}-W_{j}\|^{2}$ where $f$ represents the maximum number of malicious clients that can be tolerated and $i\to j$ denotes that $W_{j}$ belongs to the nf2n-f-2 closest updates to $W_{i}$. Subsequently, the model with the lowest score or the average of several models with the lowest scores was chosen as the global model.
  • FLTrust[8]. The method requires the server to maintain a small clean dataset, which is used to train a trusted model update $g_{0}$. A trust score for each local model update $g_{i}$ is then computed by evaluating the cosine similarity between $g_{i}$ and $g_{0}$, followed by applying a ReLU operation. The score is used as the aggregation weight of the local model in the aggregation process.
  • AGR in FheFL[27]. In this schemes, the aggregation methods use the Euclidean distance between the local model $W^{t}_{i}$ and the global model $W^{t-1}$ of the previous round as a discriminator of the quality of the local model $W^{t}_{i}$.
  • AGR in BSR-FL[31]. This aggregation method constructs a qualified aggregation model using secure cosine similarity and an incentive-based Byzantine robustness strategy. In BSR-FL, the aggragated benign model from the (t1)(t-1)-th iteration $W_{base}^{t-1}$, serves as the baseline model. The trust score $cs^{t}_{i}$ for each local model $W_{i}^{t}$ is computed as the cosine similarity between $W_{i}^{t}$ and the baseline $W_{base}^{t-1}$, and then transformed using a parameterized sigmoid function centered at 0.5. However, the BSR does not verify the length of local models, leaving the potential for spoofing.
Existing aggregation rules are either not robust enough to resist adaptive attacks[8], or too complex to be applicable for fast verification of zero-knowledge proofs. Therefore, we propose a new aggregation rule that is applicable to zero-knowledge proofs and provides an effective defense against attacks such as adaptive attacks.
Privacy-Preserving FL against Malicious Clients. Although client data is not directly exchanged in the federated learning framework, the transmission of local models still introduces potential privacy risks. In order to make PPFL resistant to the behavior of malicious participants, some works have been proposed to implement the above defense strategies in PPFL. The relevant schemes can be categorized into 3 groups according to privacy techniques. A comparison of these previous schemes with our method is summarized in Table I.
    Differential Privacy Based FL. The Biscotti[28] combined differential privacy, Shamir secret shares (SS) and Multi-Krum aggregation rule to realize a private and secure federated learning scheme based on the blockchain. SecProbe[32] aggregated local models based on the computed utility scores and applies a functional mechanism to perturb the objective function to achieve privacy. Miao et al.[24] proposed an efficient and secure federated learning (ESFL) framework based on DP. Since the nature of differential privacy techniques is to hide the sensitive information of an individual by introducing random noise in the data or computational results, it cannot avoid affecting the accuracy of the model.
    Homomorphic Encryption Based FL. SecureFL[19] implemented FLTrust construction under encryption based on packed linear homomorphic encryption. However, in this scheme, the client distributes the local gradient to two servers through secret sharing, and the privacy protection of the gradient must rely on two non-colluding servers. AegisFL[13] improved the computational efficiency of multiple aggregation rules under ciphertexts by optimizing the homomorphic encryption scheme, but it also fails to eliminate dependency on two non-colluding servers. The RVPFL[22] scheme combined homomorphic encryption and Diffie-Hellman key agreement to achieve privacy protection and robustness detection, but relies on additional participants. FheFL[27] built the scheme on homomorphic encryption in a single server. However, a trusted key distribution process is necessary to ensure the secure generation of pairwise keys among users. The scheme does not validate the client behavior, and the malicious clients can easily destroy the decryption through wrong keys. Such schemes often use homomorphic encryption for privacy, computing aggregation on ciphertexts. Decryption typically requires a trusted party or two non-colluding servers. Shared decryption keys risk exposing honest clients’ data to key holders through collusion or eavesdropping.
    Functional Encryption Based FL. Chang et al.[12] constructed a privacy-preserving federated learning scheme FEFL via functional encryption. Qian et al.[26] designed another privacy-preserving federated learning scheme by combining functional encryption with multiple cryptographic primitives. However, both schemes do not address the detection of malicious clients. The BSR-FL scheme [31] combined functional encryption and blockchain technology to construct an efficient Byzantine robust privacy-preserving federated learning framework. However, the scheme still requires two non-colluding entities (the task publisher and the server) and does not validate the magnitude of the local models submitted by clients.
III PROBLEM STATEMENT
III-A System Model
The VFEFL architecture proposed in this paper is applicable to a wide range of privacy-preserving federated learning settings and does not require any trusted third party. Its architecture involves two main roles: a server $\mathcal{S}$ and $n$ clients $\{\mathcal{C}_{i}\}_{i\in[n]}$.
Server: The server, equipped with a clean dataset $D_{0}$ and substantial computational resources, is responsible for detecting maliciously encrypted local models submitted by clients, aggregating the models from clients, and returning the aggregated model to the clients for further training. This process is iteratively performed until a satisfactory global model is achieved.
Clients: Clients have their own datasets and a certain amount of computational power, and it is expected that multiple clients will cooperate to train the model together without revealing their own data and model information.
The notations in the paper are described in Table II
TABLE II: Notations and Descriptions
Notations Descriptions
$n$ the number of clients
$m$ the dimension of model
$t$ training epoch
lrlr local learning rate
$R_{l}$ the number of local iterations
$\ell$ the label in the CC-DVFE
$W^{t}$ global model
$W_{0}$ initial global model
$W_{i}^{t}$ local model
$D_{0}$ root dataset
$D_{i}$ local dataset
$g,w_{i},u,v$ the generators of $\mathbb{G}_{1}$
$h,\hat{v}_{1},\hat{v}_{2}$ the generators of $\mathbb{G}_{2}$
Note: The superscripts of $W_{i}^{t},\vec{x}_{i}^{t},D_{i}$ denote interactions and the subscripts denote client identifiers.
III-B Design Goals
VFEFL is designed to provide the following performance and security guarantees:
  1. -
    Privacy. VFEFL should ensure that the server or channel eavesdroppers do not have access to the client’s local model to reconstruct the client’s private data.
  2. -
    Byzantine Robustness. VFEFL can resist Byzantine attacks such as Gaussian attack[15], the adaptive attack[8] and Label Flipping attack, and obtain a high-quality global model.
  3. -
    Self-contained. VFEFL can be deployed in the basic federated learning framework, i.e., without relying on two non-colluding servers and additional third parties, thus gaining more general applicability.
  4. -
    Verifiability. VFEFL ensures clients’ outputs comply with the protocol, including ciphertexts and functional key shares, preventing malicious clients from framing honest ones or disrupting aggregation. It also ensures system continuity despite partial data corruption during transmission.
  5. -
    Fidelity. VFEFL architecture causes no loss of accuracy in the absence of attacks
We will proof the above properties in Section VII.
III-C Security Model
[…]
where $\stackrel{{\scriptstyle c}}{{\approx}}$ denotes computational indistinguishability, $\lambda$ represents the security parameter, and $\hat{x}$ includes all inputs.
In the design goal of VFEFL, privacy is defined that neither the server nor eavesdroppers can obtain the local models of clients, thus preventing them from inferring the clients’ data samples during interactions between the server and clients. In the execution of VFEFL protocols, an attacker may compromise either some clients or the server, thereby gaining access to encrypted data and intermediate results. With them, the attacker seeks to extract private information. Thus, we define the ideal function of VFEFL as $f_{\mathrm{VFEFL}}$. The privacy of VFEFL can be formally described as:
[…]
where $\hat{y}$ denotes the $y$ and intermediate results, $\hat{x}_{\mathcal{S}},\hat{x}_{\mathcal{C}}$ denote the inputs of $\mathcal{S}$ and $\mathcal{C}$, respectively. The formal proof as discussed in Section VII-A.
III-D Threat Model
We consider the honest-but-curious server, which executes the protocol faithfully but may leverage techniques such as gradient inversion attacks on the client’s model to infer private data. In this paper, we categorize the clients $\mathcal{C}$ into two groups: honest clients $\mathcal{HC}$ and malicious clients $\mathcal{CC}$. An honest client honestly trains a local model based on its real dataset in a training iteration, encrypts and proofs the model, generates the correct functional key share and the proofs, and submits them to $\mathcal{S}$. A malicious client may submit random or carefully crafted model parameters to degrade the accuracy of the global model, or submit incorrect encrypted models or functional key shares to cause decryption failure. Additionally, a malicious client may eavesdrop on an honest client’s secrets by colluding with others.
Our system requires at least two clients to be honest to ensure security, even in the presence of collusion.
IV PRELIMINARIES
IV-A Groups
Prime Order Group. We use a prime-order group GGen, a probabilistic polynomial-time (PPT) algorithm that, given a security parameter $\lambda$, outputs a description $\mathcal{G}=(\mathbb{G},p,g)$, where $\mathbb{G}$ is an additive cyclic group of order $p$, $p$ is a $2\lambda$-bit prime, and $g$ is the generator of $\mathbb{G}$. For any $a\in\mathbb{Z}_{p}$, given $g^{a}$, it is computationally hard to determine $a$, a problem known as the discrete logarithm problem. For any $\vec{a}=\{a_{1},\cdots,a_{n}\}\in\mathbb{Z}^{n}_{p}$, $\vec{g}=\{g_{1},\cdots,g_{n}\}\in\mathbb{G}^{n}$, we denote $\vec{g}^{\vec{a}}=\prod_{i\in[n]}g_{i}^{a_{i}}$.
Pairing group. We use a pairing group generator PGGen, a PPT algorithm that on input the security parameter $\lambda$ returns a a description $\mathcal{PG}=\left(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e\right)$ of asymmetric pairing groups where $\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T}$ are additive cyclic groups of order $p$ for a $2\lambda$-bit prime $p$, $g$ and $h$ are generators of $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$, respectively, and $e:\mathbb{G}_{1}\times\mathbb{G}_{2}\rightarrow\mathbb{G}_{T}$ is an efficiently computable (non-degenerate) bilinear group elements. For $a,b\in\mathbb{Z}_{p}$, given $g^{a},h^{b}$, one can efficiently compute $e(g,h)^{ab}$ using the pairing $e$.
[…]
IV-B Zero Knowledge Proof
For a polynomial-time decidable relation $\mathcal{R}$, we call $w$ a witness for a statement $u$ if $\mathcal{R}(u;w)=1$. Let $L=\{u|\exists w:\mathcal{R}(u;w)=1\}$ is a language associated with $\mathcal{R}$. A zero-knowledge proof for $L$ is a protocol between prover $\mathcal{P}$ and verifier $\mathcal{V}$ where $\mathcal{P}$ convinces $\mathcal{V}$ that a common input $u\in L$ without revealing information about a witness $w$. A zero-knowledge proof should satisfy the following three properties: soundness, completeness and perfect honest-verifier zero-knowledge. Soundness refers to the probability of $\mathcal{P}$ convincing $\mathcal{V}$ to accept $u\in L$ is negligible when $u\notin L$. Completeness means that if $u\in L$, $\mathcal{V}$ accept it with non-negligible probability. And the zero-knowledge implies that $\mathcal{V}$ cannot revealing the information of $w$. We denote by
$\mathsf{PoK}\{(w):\varphi(w)\}$
a proof of knowledge of a secret value $w$ such that the public predicate $\varphi(w)$ holds.
Notably, an interactive proof-of-knowledge protocol can be transformed into a non-interactive proof-of-knowledge through the application of the Fiat-Shamir heuristic [16].
IV-C Cross-Ciphertext Decentralized Verifiable Functional Encryption
Verifiable functional encryption (VFE) [2] was introduced to capture ciphertexts and decryption keys’ correctness under fine-grained control. Informally, VFE requires that for any ciphertext $C$ that passes a public verification procedure, there exists a unique message $m$ such that, for any valid function description $f$ and any corresponding function key $dk_{f}$ that also pass public verification, the decryption algorithm outputs $f(m)$ when given $C$, $dk_{f}$, and $f$ as inputs.
Nguyen et al.[25] proposed a decentralized multi-client verifiable functional encryption scheme. The scheme focuses on verifying properties of individual ciphertexts. In our work, we propose a verifiable functional encryption scheme that enables cross-ciphertext verification, allowing the validation of specific relations among multiple ciphertexts in various application scenarios. Next, we give the formal definition of our Cross-Ciphertext Decentralized Verifiable multi-client Functional Encryption for inner product (CC-DVFE).
Definition 1
A Cross-Ciphertext Decentralized Verifiable Multi-Client Functional Encryption for Inner Product among a set of $n$ clients $\left(\mathcal{C}_{i}\right)_{i\in[n]}$ with a $m$-dimensional vector on $\mathcal{M}^{m}$ contains eight algorithms.
[…]
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: Each client $\mathcal{C}_{i}$ generates own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs the verifiation key for ciphertexts $vk_{CT}$ and for functional key share $vk_{Dk}$, and the public key $pk$.
[…]
Correctness. Given any set of message $\left(\vec{x}_{1},\cdots,\vec{x}_{n}\right)\in\mathcal{M}^{n\times m}$ and the auxiliary information auxaux and any function vector $\vec{y}\in\mathbb{Z}_{p}^{n}$, We say that an CC-DVFE scheme is correct if and only if following equation holds.
$\begin{split}\Pr\left[\begin{array}[]{l|l}&\texttt{SetUp}(\lambda)\rightarrow\mathrm{pp}\\ &\texttt{KeyGen}(pp)\rightarrow((ek_{i},sk_{i})_{i\in[n]},\\ &\quad vk_{CT},vk_{DK},pk),\\ \texttt{DKeyComb}(&\forall i\in[n],\ \texttt{Encrypt}(ek_{i},\vec{x}_{i},\\ \quad({dk}_{y,i})_{i\in[n]},\ell_{y},&\quad aux,\ell,(\ell_{Enc,j})_{j\in[m]},pp)\\ \quad pk)\to dk&\quad\rightarrow(C_{\ell,i},\pi_{CT,i})\\ \texttt{Decrypt}((C_{\ell,i})_{i\in[n]},&\forall i\in[n],\ \texttt{DKeyGenShare}(\\ \quad{dk},\vec{y})\to&\quad sk_{i},pk,y_{i},\ell_{y},pp)\\ \quad\left(\langle\vec{x}^{\top},\vec{y}\rangle\right)_{j\in[m]}&\quad\rightarrow(dk_{y,i},\pi_{\mathrm{DK},i})\\ &\texttt{VerifyCT}((C_{\ell,i})_{i\in[n]},\pi_{CT},\\ &\quad vk_{CT})\to 1\\ &\texttt{VerifyDK}(({dk}_{y,i})_{i\in[n]},\pi_{DK},\\ &\quad vk_{DK},pp)\to 1\\ \end{array}\right]=1.\end{split}$
IV-D Our New Aggregation Rule
Inspired by the aggregation rule of FLTrust[8], we similarly introduce root dataset $D_{0}$ to bootstrap trust to achieve Byzantine robustness against malicious clients and consider both the direction and magnitudes of the local models of clients. However, FLTrust contains cosine similarity operations that are difficult to complete the verification using a simple zero-knowledge proof, which is avoided in our new aggregation rule.
A malicious client can arbitrarily manipulate the direction and magnitudes of the local model to degrade the accuracy of the global model, so our robust aggregation rule restricts both ways. First, we assume that the server has a small clean dataset $D_{0}$ and train on it to get a baseline model $W^{t}_{0}$. Without considering the magnitudes of the model, we use the inner product of the local client model $W^{t}_{i}$ and the baseline model $W^{t}_{0}$ to measure the directional similarity between them. In order to avoid malicious users to enhance the attack effect of malicious vectors by amplifying the gradient, we incorporate the inner product value of the local client model $W^{t}_{i}$ itself into the robustness aggregation rule. We further consider that when the inner product value is negative, the model still has some negative impact on the global model of the aggregation, so we use the Relu function to further process the inner product value. The Relu function is defined as follows:
$\text{ReLu}\left(x\right)=\begin{cases}x,&\text{if }x>0\\ 0,&\text{others}\end{cases}.$
We define the robust aggregation rule as follows:
$W^{*}=\sum_{i\in[n]}\text{Relu}\left(\frac{\langle W^{t}_{i},W^{t}_{0}\rangle}{\langle W^{t}_{i},W^{t}_{i}\rangle}\right)W^{t}_{i}$
At the same time, in order to avoid the global model unable to statute to a good level of accuracy, we normalize each epoch of global model to a certain magnitude. We use the $W_{0}^{t}$ as a reference to normalize each epoch of global model. That is, theglobal gradient $W^{t}$ of for the $t$-th iteration is defined as follows:
[…]
Denote $y_{i}=\text{Relu}\left(\frac{\langle W^{t}_{i},W^{t}_{0}\rangle}{\langle W^{t}_{i},W^{t}_{i}\rangle}\right)$ and $W^{*}$ can be obtained from the functional encryption for inner product. The robustness of this aggregation rule is proved in Section VII.
V OUR DECENTRALIZED VERIABLE FUNCTIONAL ENCRYPTION
V-A Construction of CC-DVFE
Figure 1: The workflow and main components of CC-DVFE
Although verifiable functional encryption schemes supporting weight aggregation for encrypting local models are now readily available, their application to privacy-preserving and robust federated learning faces the following two challenges. The first one is that the current ciphertext verification in verifiable functional encryption supports range proofs for individual elements but cannot verify Byzantine robustness rules. Secondly, the construction of zero-knowledge proofs for the Byzantine robust aggregation rule faces compatibility issues: existing inner-product arguments [7] and functional encryption schemes cannot be directly combined for verification. Specifically, the current inner-product argument fails to achieve the zero-knowledge property, and the ciphertext structure in functional encryption does not support direct verification of Byzantine robust aggregation rules.
To address these limitations, we propose a novel verifiable functional encryption scheme for the new robust aggregation rule. Next we describe the complete DVFE scheme construction, with eight algorithms. The flowchart of the algorithm DVFE is shown in Figure 1.
  • SetUp($\lambda$) $\to pp$: Takes as input the security parameter $\lambda$ and outputs the public parameters $pp=(\mathcal{PG},\mathcal{DG},h_{p},\mathcal{H}_{1},\mathcal{H}_{2},$ $\mathcal{H}^{\prime}_{1},\mathcal{H},\ell_{D})$, where $\mathrm{\texttt{PGGen}}(\lambda)\to\mathcal{PG}=(\mathbb{G}_{1},\mathbb{G}_{2},\mathbb{G}_{T},p,g,h,e)$ is a pairing group, $g$ the generator of $\mathbb{G}_{1}$, $h$ the generator of $\mathbb{G}_{2}$, GenClassGroup $(\lambda,p)\to\mathcal{DG}=(\tilde{s},f,\hat{h}_{p},$ $\hat{G},F,\hat{G}^{p})$ is a DDH group with an easy DL subgroup, $\hat{t}\leftarrow\mathcal{D}_{p}$ is sampled, and $h_{p}=\hat{h}_{p}^{\hat{t}}$ is set. Four full-domain hash functions are defined as follows:
    $\mathcal{H}_{1}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{1}^{2},$ $\mathcal{H}_{2}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{2}^{2},$
    $\mathcal{H}^{\prime}_{1}$ $:\{0,1\}^{*}\rightarrow\mathbb{G}_{1},$ $\mathcal{H}$ $:\{0,1\}^{*}\rightarrow\mathbb{Z}_{p}.$
    And $\ell_{D}=(\{0,1\}^{*})$ is the initialization label.
  • KeyGen(pppp) $\to\left((ek_{i},sk_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$: It is executed by each client $\mathcal{C}_{i}$ to generate own encryption key $ek_{i}$ and secret key $sk_{i}$, and eventually outputs a verifiation key for ciphertexts $vk_{CT}$ and functional key share $vk_{Dk}$, and a public key $pk$. A client selects $\hat{k}_{i}=(\hat{k}_{i,1},\hat{k}_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, $t_{i}=(t_{i,1},t_{i,2})\overset{\mathdollar}{\leftarrow}\mathcal{D}^{2}_{p}$ and $s_{i}=(s_{i,1},s_{i,2})\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2}$, then computes $T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}$ and
    $d_{i}=\left(f^{\hat{k}_{i,b}}\cdot\left(\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\prod_{i>i^{*}}^{n}T_{i^{*},b}^{-1}\right)^{t_{i,b}}\right)_{b\in[2]}$
    and commits $s_{i}$ as
    $\mathrm{com}_{i}=(v)^{s_{i}}$
    where $v=\mathcal{H}_{1}(\ell_{D})\in\mathbb{G}_{1}^{2}$. For each client $\mathcal{C}_{i}$, encryption key is $ek_{i}=s_{i}$, secret key is $sk_{i}=(s_{i},\hat{k}_{i},t_{i})$ , and public key is $pk=\left(T_{i},\ d_{i}\right)_{i\in[n]}$. The verification key for ciphertexts is $vk_{CT}=\left(\mathrm{com}_{i}\right)_{i\in[n]}$, while for functional keys it is $vk_{DK}=\left(T_{i},d_{i},\mathrm{com}_{i}\right)_{i\in[n]}$.
[…]
    $\text{PoK}\left\{(s_{i},\vec{x}_{i}):\begin{aligned} V_{i}=(\bar{u})^{s_{i}}\cdot\vec{w}^{\vec{x}_{i}}\wedge\\ y_{i}=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}.\end{aligned}\right\}$
    where $\bar{u}=\prod_{j}^{m}u_{j}$, $u_{j}=\mathcal{H}_{1}(\ell||j)$, $\vec{w}^{\vec{x}_{i}}=\prod_{j\in[m]}(w_{j})^{x_{i,j}}$. Outputs ciphertext $C_{\ell,i}$ and proof $\pi_{CT,i}$.
[…]
    $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{s_{i,b}\cdot y_{i}}\right)_{b\in[2]}$
    where $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b})\in\mathbb{G}_{2}^{2},b\in[2]$. It re-computes $\mathrm{com}_{i}$ and generates a proof $\pi_{DK,i}$ for the relation $\mathcal{R}_{DKeyGenShare,i}$ on input $\left(pk,com,{dk}_{y,i},\ell_{y}\right)$. The $\pi_{DK,i}$ is described as
    $\text{PoK}\left\{(s_{i},t_{i},\hat{k}):\begin{aligned} T_{i}=h_{p}^{t_{i}}\wedge\\ d_{i}=(f^{\hat{k}_{b}}K_{\sum,i,b}^{t_{i,b}})_{b\in[2]}\wedge\\ dk_{i}=\left((\hat{v}_{b})^{\hat{k}}\cdot h^{s_{i,b}\cdot y_{i}}\right)_{b\in[2]}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}\end{aligned}\right\}$
    where $K_{\sum,i,b}=\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\in G$, $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b}),b\in[2],v=\mathcal{H}_{1}(\ell_{D})$. Outputs a functional decryption key share ${dk}_{y,i}$ and $\pi_{DK,i}$.
[…]
    $h^{dk_{y}}=\left(\frac{\prod_{i\in[n]}dk_{i,b}}{(\hat{v}_{b})^{d}}\right)_{b\in[2]}$
    where $\hat{v}_{b}=\left(\mathcal{H}_{2}(\ell_{y,b})\right),b\in[2]$. Outputs the functional decryption key $dk=h^{dk_{y}}$.
[…]
    $E_{j}^{\langle\vec{x}^{\top},\vec{y}\rangle}=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{j,1},h^{dk_{y,1}})e(u_{j,2},h^{dk_{y,2}})}$
    and solves the discrete logarithm in basis $E_{j}$ to return $\langle\vec{x}^{\top},\vec{y}\rangle$, where $\vec{x}^{\top}=(x_{1,j},x_{2,j},\dots,x_{n,j})\in\mathcal{M}^{n}$ denotes the vector consisting of the $j$-th component of each of $n$ clients. It outputs $(\langle\vec{x}^{\top},\vec{y}\rangle)_{j\in[m]}$ or $\bot$.
Correctness.We have
$f^{\sum_{i\in[n]}\hat{k}_{i}}$ $=\prod_{i\in[n]}d_{i}$
$d$ $=\sum_{i\in[n]}\hat{k}_{i}=\log_{f}f^{\sum_{i\in[n]}\hat{k}_{i}}$
$h^{dk_{y}}$ $=\left(\frac{\prod_{i\in[n]}dk_{i,b}}{(\hat{v}_{b})^{d}}\right)_{b\in[2]}$
$=\left(\frac{\prod_{i\in[n]}(\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{s_{i,b}\cdot y_{i}}}{(\hat{v}_{b})^{\sum_{i\in[n]}\hat{k}_{i}}}\right)_{b\in[2]}$
$=\left(h^{\sum_{i\in[n]}s_{i,b}\cdot y_{i}}\right)_{b\in[2]}$
$(E_{j})^{\langle\vec{x}^{\top},y\rangle}$ $=\frac{\prod_{i\in[n]}e(c_{l,i,j},h^{y_{i}})}{e(u_{j,1},h^{dk_{y,1}})e(u_{j,2},h^{dk_{y,2}})}$
$=\frac{\prod_{i\in[n]}e(u^{s_{i}}\cdot(w_{j})^{x_{i,j}},h^{y_{i}})}{e(u_{j,1},h^{dk_{y,1}})e(u_{j,2},h^{dk_{y,2}})}$
$=e(w_{j},h)^{\langle\vec{x}^{\top},y\rangle}$
where $\vec{x}^{\top}=(x_{1,j},x_{2,j},\dots,x_{n,j})\in\mathcal{M}^{n}$ denote the vector composed of the $j$-th parameter of the local models from each of the $n$ clients, and then solves the discrete logarithm in basis $E_{j}$ to return $\langle\vec{x}^{\top},y\rangle$. Performing $m$ decryption operations on $\{\left\{c_{\ell,i,j}\right\}_{i\in[n]}\}_{j\in[m]}$ returns the result $\{\langle\vec{x}^{\top},y\rangle\}_{j\in[m]}$.
Theorem 1
The decentralized verifiable multi-client functional encryption for inner product is static-IND-secure under the DDH, multi-DDH and HSM assumptions, as Definition 6. More precisely, we have
$\mathrm{Adv_{CC-DVFE}^{sta-ind}}(t,q_{E},q_{K})\leq\$$q_{E}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$
$+q_{K}\mathrm{Adv}^{\mathrm{zk}}_{\prod_{DKeyGenShare}}(t)$
$+\mathrm{Adv^{std-ind}_{LDSUM}}(t,q_{K})$
$+2q_{E}(2\mathrm{Adv}^{ddh}_{\mathbb{G}_{1}}(t)+\frac{1}{p})$
$+2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t+4q_{E}\times t_{\mathbb{G}_{1}})$
where
    $\mathrm{Adv_{CC-DVFE}^{sta-ind}}(t,q_{E},q_{K})$ is the best advantage of any PPT adversary running in time $t$ with $q_{E}$ encryption queries and $q_{K}$ key share queries against the static-IND-secure game of the CC-DVFE scheme;
    $\mathrm{Adv}^{\mathrm{zk}}_{\prod_{Encrypt}}(t)$ is the best advantage of any PPT adversary running in time $t$ against the zero-knowledge property of the $\prod_{Encrypt}$ scheme;
    $\mathrm{Adv}^{\mathrm{zk}}_{\prod_{DKeyGenShare}}(t)$ is the best advantage of any PPT adversary running in time $t$ against the zero-knowledge property of the $\prod_{DKeyGenShare}$ scheme;
    $\mathrm{Adv^{std-ind}_{LDSUM}}(t,q_{K})$ is the best advantage of any PPT adversary running in time $t$ with $q_{K}$ encryption queries against the IND-security game of the LDSUM scheme[25].
    $\mathrm{Adv}^{ddh}_{\mathbb{G}_{1}}(t)$ is the best advantage of breaking the DDH assumption in $\mathbb{G}_{1}$ within time $t$.
Proof 1
It is worth noting that, to achieve a non-negligible advantage in winning the game, the adversary $\mathcal{A}$ must leave at least two honest clients, each possessing two distinct messages $(X_{0},X_{1})$ for encryption queries. We define games as follows:
Game $G_{0}$: The challenger plays as this game described in the Definition 6, with the set of $\mathcal{CC}$ of the corrupted clients are established before initialization. Note that the hash function $\mathcal{H}_{1},\mathcal{H}_{2},\mathcal{H}^{\prime}_{1},\mathcal{H}$ is modeled as a random oracle RO onto $\mathbb{G}_{1}^{2},\mathbb{G}_{2}^{2},\mathbb{G}_{1},\mathbb{Z}_{p}$.
Game $G_{1}$: This game is as $G_{0}$, except that for generating proofs: $\pi_{CT}$ and $\pi_{DK}$ in the QEncrypt and QDKeyGenShare.
    $\pi_{\mathrm{CT}}$ is produced by a simulator $\mathrm{Sim}_{\mathrm{Encrypt}}$ on the input $(\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]},C_{\ell,i},\vec{x}_{0},\mathrm{com}_{i})$ if $\mathcal{C}_{i}$ is queried in QEncrypt for the input $(i,X^{0},X^{1}\in[0,2^{q}-1]^{n\times m},\ell,\left\{\ell_{\mathrm{Enc}}\right\}_{j\in[m]})$.
[…]
Given $q_{E}$ encryption queries and $q_{K}$ functional key share queries, with the zero-knowledge property of NIZK proofs, we have
$\left|\mathrm{Adv}_{G_{0}}-\mathrm{Adv}_{G_{1}}\right|$
$\leq$ $\ q_{E}\mathrm{Adv^{zk}_{\prod_{Encrypt}}}(t)+q_{K}\mathrm{Adv^{zk}_{\prod_{DKeyGenShare}}}(t)$
Game $G_{2}$: This game is as $G_{1}$, except for functional key share queries $\mathrm{QDKeyGen}(i,y)$:
    if $i$ is a corrupted index, then the challenger computes
    $dk_{i,j}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{s_{i,b}\cdot y_{i}}\right)_{b\in[2]}$
    if $i$ is the last non-corrupted index, the challenger computes
    $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{\sum_{t\in\mathcal{HC}}s_{t,1}\cdot y_{t}}\right)_{b\in[2]}$
    for other non-corrupted index $i$, the challenger computes
    $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{0},\right)_{b\in[2]}$
Given $q_{K}$ key share queries, by the (static) IND-security of the LDSUM scheme in [25], we have $\left|\mathrm{Adv}_{G_{1}}-\mathrm{Adv}_{G_{2}}\right|\leq\mathrm{Adv^{std-ind}_{LDSUM}}(t,q_{K})$ under the HSM assumptions.
Game $G_{3}$: We simulate the answers to $\mathcal{H}_{1}$ to any new RO-query of $\mathbb{G}_{1}^{2}$ by a truly random pair in $\mathbb{G}_{1}^{2}$ on the fly. Namely, $(g^{a},g^{b})$, $a,b,\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. The simulation remains perfect, and so $\mathrm{Adv}_{2}=\mathrm{Adv}_{3}$.
Game $G_{4}$: We simulate the answers to $\mathcal{H}^{\prime}_{1}$ to any new RO-query of $\mathbb{G}_{1}$ by a truly random element of $w_{j}=g^{\hat{w}_{j}}$, with $\hat{w}_{j}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. $\mathrm{Adv}_{3}=\mathrm{Adv}_{4}$.
Game $G_{5}$: We simulate the answers to $\mathcal{H}_{1}$ to any new RO-query of $\mathbb{G}^{2}_{1}$ by a truly random pair in the span of $g^{\mathbf{a}}$ for $\mathbf{a}:=\left(1,\ a\right)^{T}$, with $a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. In other words, $u=g^{\hat{u}}=g^{\mathbf{a}r_{\ell}}=\left(g^{r_{\ell}},g^{ar_{\ell}}\right),r_{\ell}\overset{\mathdollar}{\leftarrow}\mathrm{RF}_{\mathbb{Z}_{p}}(\ell)$. From the following Lemma 1, $\left|\mathrm{Adv}_{4}-\mathrm{Adv}_{5}\right|\leq 2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t+4q_{E}\times t_{\mathbb{G}_{1}})$.
Game $G_{6}$: We simulate any QEncrypt query as the encryption of $X_{i}^{0}$ instead of $X_{i}^{b}$ and go back for the answers to $\mathcal{H}_{1}$ to any new RO-query by a truly random pair in $\mathbb{G}^{2}_{1}$ and to $\mathcal{H}^{\prime}_{1}$ to a truly random element in $\mathbb{G}_{1}$. From the following Lemma 2, we have $\mathrm{Adv}_{5}-\mathrm{Adv}_{6}\leq 2q_{E}(2\mathrm{Adv}^{ddh}_{\mathbb{G}}(t)+\frac{1}{p})$.
In addition, it is clear that in the Game $G_{6}$ the advantage of any adversary is exactly 0 since $b$ does not appear anywhere. In other words, $\mathrm{Adv}_{6}=0$, which concludes the proof.
Lemma 1
If the Multi-DDH assumption holds, then no polynomial-time adversary can distinguish between Game $G_{4}$ and Game $G_{5}$ with a non-negligible advantage.
Proof 2
Given a multi-DDH instance
$\left((g^{x},(g^{y_{q}},g^{z_{q}})_{q\in[q_{E}]})\mid x,y_{q}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\right),$
the challenger flips an unbiased coin with $\{0,1\}$ and obtains a bit $\mu\in\{0,1\}$. if $\mu=0,z_{q}=xy_{q}$; if $\mu=1,z_{q}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. Suppose there exists an PPT adversary $\mathcal{A}$ who can distinguishes
$A=\left\{g^{a},(g^{r_{q}},g^{ar_{q}})_{q\in[q_{E}]}\mid r_{q}\overset{\mathdollar}{\leftarrow}\mathrm{RF}_{\mathbb{Z}_{p}}(\ell_{q}),a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},q\in[q_{E}]\right\}$
and
$B=\left\{g^{a},(g^{b},g^{c})_{q\in[q_{E}]}\mid a,b,c\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},q\in[q_{E}]\right\}$
by a advantage $\mathrm{Adv}(t)$, where $\mathrm{Adv}(t)$ denotes the advantage of an adversary in distinguishing $G_{4}$ from $G_{5}$ within time $t$. There exists a simulator $\mathcal{B}$ who can use $\mathcal{A}$ to break the multi-DDH assumption as follows. $\mathcal{B}$ sets $C=(g^{x},(g^{y_{q}},g^{z_{q}})_{q\in[q_{E}]})$. $\mathcal{B}$ sends $C$ to $\mathcal{A}$ and make the adversary guess that C is equal to A or B. If $\mathcal{A}$ the believes that $C=A$ outputs $\omega=0$, $C=B$ outputs $\omega=1$. $\mathcal{B}$ outputs the its guess $\mu^{\prime}$ on $\mu$: $\mu^{\prime}=\omega$. It follows from Definition 5 that the total advantage with which $\mathcal{B}$ can break the multi-DDH assumption in $\mathbb{G}_{1}$ is $\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t+4q_{E}\times t_{\mathbb{G}_{1}})$. Therefore,
$\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t+4q_{E}\times t_{\mathbb{G}_{1}})$
$=$ $|\operatorname{Pr}\left[\mu^{\prime}=\mu\mid\mu=0\right]\times\operatorname{Pr}[\mu=0]$
$-\operatorname{Pr}\left[\mu^{\prime}=\mu\mid\mu=1\right]\times\operatorname{Pr}[\mu=1]|$
$\geq$ $(\frac{1}{2}+\mathrm{Adv}(t))\times\frac{1}{2}-\frac{1}{2}\times\frac{1}{2}$
$=$ $\frac{1}{2}\mathrm{Adv}(t)$
Therefore, $\left|\mathrm{Adv}_{4}-\mathrm{Adv}_{5}\right|\leq 2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t+4q_{E}\times t_{\mathbb{G}_{1}})$.
Lemma 2
If the DDH assumption holds, then no polynomial-time adversary can distinguish between Game $G_{5}$ and Game $G_{6}$ with a non-negligible advantage.
Proof 3
The gap between $G_{5}$ and $G_{6}$ will be proven using a hybrid technique on the RO-queries. Accordingly, we index the following games by $q$, where $q=1,\cdots,q_{E}$.
  • $G_{5.1.1}$: This is exactly game $G_{5}$. Thus, $\mathrm{Adv}_{5}=\mathrm{Adv}_{5.1.1}$
[…]
  • $G_{5.q.3}:$ We now change the ciphertext response to encryption queries, shifting $\{c_{\ell,i,j}\}_{j\in[m]}=\{(u_{j})^{s_{i}}\cdot(w_{j})^{X^{b}_{i,j}}\}_{j\in[m]}$ to $\{c_{\ell,i,j}\}_{j\in[m]}=\{(u_{j})^{s_{i}}\cdot(w_{j})^{X^{0}_{i,j}}\}_{j\in[m]}$ where $u_{j}$ corresponds to the $q$-th RO-query of $\mathbb{G}^{2}_{1}$ and $w_{j}$ corresponds to the $q$-th RO-query of $\mathbb{G}_{1}$. From the following Lemmas 4, we can get $\mathrm{Adv}_{5.q.2}=\mathrm{Adv}_{5.q.3}$.
  • $G_{5.q+1.1}:$ This transition is the reverse of that $G_{5.q.1}$ to $G_{5.q.2}$, namely, we switch back the distribution of $u$ computed on the $q$-th RO-query from uniformly random over $\mathbb{G}^{2}$ (conditioned on the fact that $\hat{u}(\hat{a})^{\bot}\neq 0$) to uniformly random in the span of $g^{\mathbf{a}}$. Thus, we can get $\left|\mathrm{Adv}_{5.q.3}-\mathrm{Adv}_{5.q+1.1}\right|\leq 2\mathrm{Adv}^{ddh}_{\mathbb{G}}(t)+\frac{1}{p}$.
Therefore, $\left|\mathrm{Adv}_{5}-\mathrm{Adv}_{6}\right|\leq 2q_{E}(2\mathrm{Adv}^{ddh}_{\mathbb{G}}(t)+\frac{1}{p})$.
Lemma 3
For any PPT adversary, let $\mathrm{Adv}(t)$ denote its advantage in distinguishing the distributions
$\mathcal{D}_{0}=\{g,g^{r},g^{ar}\mid a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},\ r=RF_{\mathbb{Z}_{p}}(\ell)\}$
and
$\mathcal{D}_{1}=\{g,g^{u_{1}-u_{2}a},g^{u_{1}a+u_{2}}\mid a,u_{1},u_{2}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\}.$
Let $\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$ denote the advantage in solving the DDH problem in $\mathbb{G}_{1}$. Then it holds that $\mathrm{Adv}(t)\leq 2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$.
Proof 4
Given a DDH instance $(g^{x},g^{y},g^{z}\mid x,y\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p})$, the challenger flips an unbiased coin with $\{0,1\}$ and obtains a bit $\mu\overset{\mathdollar}{\leftarrow}\{0,1\}$. if $\mu=0,z=xy$; if $\mu=1,z\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$. Suppose there exists an PPT adversary $\mathcal{A}$ who can distinguishes $A=\{g,g^{r},g^{ar}\mid a\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p},r=RF_{\mathbb{Z}_{p}}(\ell)\}$ and $B=\{g,g^{u_{1}-u_{2}a},g^{u_{1}a+u_{2}}\mid a,u_{1},u_{2}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\}$ within time $t$ by a advantage $\mathrm{Adv}(t)$, there exists a simulator $\mathcal{B}$ who can use $\mathcal{A}$ to break the DDH assumption as follows.
$\mathcal{B}$ sets $C=(X,Y,Z)$. $\mathcal{B}$ sends $C$ to $\mathcal{A}$ and make the adversary guess that $C$ is equal to $A$ or $B$. If $\mathcal{A}$ believes that $C=A$ outputs $\omega=0$; $C=B$ outputs $\omega=1$. $\mathcal{B}$ outputs $\mu^{\prime}=\omega$. Let $\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$ denote the best advantage of breaking the DDH assumption in $\mathbb{G}_{1}$ within time $t$. Therefore,
$\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$
$=$ $\left|\operatorname{Pr}\left[\mu^{\prime}=\mu\middle|\mu=0\right]\operatorname{Pr}[\mu=0]-\operatorname{Pr}\left[\mu^{\prime}=\mu\middle|\mu=1\right]\operatorname{Pr}[\mu=1]\right|$
$\geq$ $(\frac{1}{2}+\mathrm{Adv}(t))\times\frac{1}{2}-\frac{1}{2}\times\frac{1}{2}$
$=$ $\frac{1}{2}\mathrm{Adv}(t)$
Therefore, $\mathrm{Adv}(t)\leq 2\mathrm{Adv}_{\mathbb{G}_{1}}^{ddh}(t)$.
Lemma 4
In the adversary’s point of view, game $G_{5.q.2}$ is the same as game $G_{5.q.3}$, thus, $\mathrm{Adv}_{5.q.2}=\mathrm{Adv}_{5.q.3}$, for $q=1,\cdots,q_{E}$.
Proof 5
We then prove this does not change the adversarys view from $G_{5.q.2}$ to $G_{5.q.3}$. Obviously, if the output of the $q$-th RO-query is not used by QEncrypt-queries, then the games $G_{5.q.2}$ and $G_{5.q.3}$ are identical. However, we can also demonstrate that this holds when the $q$-th RO-queries are genuinely involved in QEncrypt-queries. The prove can be divided into two step.
    Step 1: We show that there exists a PPT adversary $\mathcal{B^{*}}$ such that, for $t=2,3$, $\mathrm{Adv}_{5.q.t}=\left(p^{2m}+1\right)^{-n}\mathrm{Adv}^{*}_{5.q.t}(\mathcal{B^{*}})$, holds. Here, the games $G^{*}_{5.q.2}$ and $G^{*}_{5.q.3}$ are selective variants of games $G_{5.q.2}$ and $G_{6.q.3}$, respectively, and the set of indices submitted to QCorrupt queries are known for the challenger before the initialization phase. Fisrtly, the adversary $\mathcal{B^{*}}$ guesses for all $i\in[n],Z_{i}\overset{\mathdollar}{\leftarrow}\mathbb{Z}^{2\times m}_{p}\cup\{\bot\}$, which it sends to its selective game $G^{*}_{5.q.t}$. Specifically, each guess $Z_{i}$ is either a pair of vector $(X_{i}^{0},X_{i}^{1})$ queried to QEncrypt, or $\bot$, indicating no query was made to QEncrypt. Then, it simulates $\mathcal{A}$’s view using its own oracles. When $\mathcal{B}^{*}$ guesses correctly, denoted by event $E$, it simulates $\mathcal{A}$’s view exactly as in $G_{5.q.t}$. If the guess is wrong, then $\mathcal{B}$ stops the simulation and outputs a random bit $\beta$. Since event $E$ happens with probability $(p^{2m}+1)^{-n}$ and is independent of the view of adversary $\mathcal{A}$, we can get that $\mathrm{Adv}^{*}_{5.q.t}(\mathcal{B^{*}})$ is equal to
    $|\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=0,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}$
    $-\operatorname{Pr}\left[G_{5.q.t}^{\star}\middle|b=1,E\right]\cdot\operatorname{Pr}[E]+\frac{\operatorname{Pr}[\neg E]}{2}|$
    $=$ $\operatorname{Pr}[E]\cdot|\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=0,E\right]-\operatorname{Pr}\left[G_{5.q\cdot t}^{\star}\middle|b=1,E\right]|$
    $=$ $\left(p^{2m}+1\right)^{-n}\cdot\operatorname{Adv}_{5.q.t}.$
    Step 2: We assume the matrix $(Z_{i})_{i\in[n]}$ sent by $\mathcal{B}^{*}$ are consistent. That is, Specifically, it does not cause the game to terminate prematurely and Finalize on $\beta$ does not return a random bit independent of $b^{\prime}$, conditioned on $E^{\prime}$. Now, we demonstrate that the games $G^{*}_{5.q.2}$ and $G^{*}_{5.q.3}$ are identically distributed. To prove this, we leverage the fact that the following two distributions are identical, for any choice of $\gamma_{j}$:
    $\left\{\left(s_{i}\right)_{i\in[n],Z_{i,j}=(X_{i,j}^{0},X_{i,j}^{1})}\right\}_{j\in[m]}$
    and
    $\left\{\left(s_{i}+(\hat{a})^{\bot}\cdot\gamma_{j}(X_{i,j}^{b}-X_{i,j}^{0})\right)_{i\in[n],Z_{i,j}=(X_{i,j}^{0},X_{i,j}^{1})}\right\}_{j\in[m]}$
    where $(\hat{a})^{\bot}:=(-a,1)^{T}\in\mathbb{Z}^{2}_{p}$ and $s_{i}\leftarrow\mathbb{Z}^{2}_{p}$, for all $i=1,\cdots,n$. They are identical because the $s_{i}$ values are independent of the $Z_{i,j}$, a property that holds due to the selective nature of our setting. This independence may not necessarily apply in the case of adaptive QEncrypt queries. Thus, we can re-write $s_{i}$ into $s_{i}+(\hat{a})^{\bot}\cdot\gamma_{j}(X_{i,j}^{b}-X_{i,j}^{0})$ without changing the distribution of the game.
    We now show the additional terms $\Delta_{i,j}=(\hat{a})^{\bot}\cdot\gamma_{j}(X_{i,j}^{b}-X_{i,j}^{0})$ manifest in the adversary’s view.
    1. 1.
      They do not appear in the output of QCorrupt.
    2. 2.
      They might appear in $\text{QDKeyGen}(i,y)$. In practice, if different clients have the same function for each dimension aggregation, only one key can be generated, and if each dimension aggregation function is different, $m$ keys can be generated. For generality, we consider the case of m keys.
        if $i$ is a corrupted index, then the challenger computes $\{dk_{i,j}\}_{j\in[m]}=\{((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{(s_{i,b}+\Delta_{i,j,b})\cdot y_{i}})_{b\in[2]}\}_{j\in[m]}.$ As $i\in\mathcal{CC},Z_{i}\neq\bot,X_{i}^{1}=X_{i}^{0}$, we can get that $\left\{\Delta_{i,j}=0^{2}\right\}_{j\in[m]}$. Therefore in this case, the decryption key remains the same upon change.
        if $i$ is the last honest index, for $j\in[m]$, the challenger computes
        $dk_{i,j}=((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{\sum_{t\in\mathcal{HC}}(s_{t,b}+\Delta_{t,j,b})\cdot y_{t}})$
        By the constraints for $E^{\prime}$, as $i\in\mathcal{HC}$, $Z_{i}\neq\bot$ and $\{\langle X^{\top 0}_{j},\vec{y}\rangle=\langle X^{\top 1}_{j},\vec{y}\rangle\}_{j\in[m]}$, we can get that, for $j\in[m]$,
        $\sum_{i\in[n]:Z_{i,j}=(X_{i,j}^{0},X_{i,j}^{1})}y_{i}(X_{i,j}^{b}-X_{i,j}^{0})=0.$
        Consider the constraints on queries to corrupt clients, as $i\in\mathcal{CC}$ and $X^{0}_{i}=X^{1}_{i}$, for $j\in[m]$,
        $\sum_{i\in\mathcal{CC}:Z_{i,j}=(X_{i,j}^{0},X_{i,j}^{1})}y_{i}(X_{i,j}^{b}-X_{i,j}^{0})=0$
        is hold, and we can get that, for $j\in[m]$,
        $\sum_{i\in\mathcal{HC}}\Delta_{i,j}\cdot y_{i}=\textbf{0}.$
        Thus, for $j\in[m]$,
        $\sum_{t\in\mathcal{HC}}(s_{t}+\Delta_{t,j})\cdot y_{t}=\sum_{t\in\mathcal{HC}}s_{t}\cdot y_{t}.$
        Therefore in this case, the decryption key remains the same upon change.
        for other honest index $i$, the challenger computes
        $dk_{i}=\left((\hat{v}_{b})^{\hat{k}_{i}}\cdot h^{0}\right)_{b\in[2]}$
        This case does not involve the encryption keys.
      So in $\text{QDKeyGen}(i,y)$, it’s indistinguishable before and after the transformation.
    3. 3.
      They appear in the output of the QEncrypt-queries which use $u_{j}$ computed on the q-th RO-query. Since for all others, the vector $u_{j}$ lies in the span of $g^{\hat{a}}$, i.e., $u_{j}=g^{\hat{u}_{j}},w_{j}=g^{\hat{w}_{j}}$ and $(\hat{a})^{\top}(\hat{a})^{\bot}=0$. We thus have, for $j\in[m]$,
      $C_{i,j}=(u_{j})^{s_{i}+(\hat{a})^{\bot}\cdot\gamma_{j}(X_{i,j}^{b}-X_{i,j}^{0})}(w_{j})^{X_{i,j}^{b}}.$
      Since $\hat{u}_{j}(\hat{a})^{\bot}\neq 0$, we choose $\gamma_{j}=\frac{-\hat{w}_{j}}{\hat{u}_{j}(\hat{a})^{\bot}}\mod{p}$ and then $C_{i,j}=(u_{j})^{s_{i}}(w_{j})^{X_{i,j}^{0}}$. We stress that $\gamma_{j}$ is independent of the index $i$, and so this simultaneously converts all the encryptions of $X_{i}^{b}$ into encryptions of $X_{i}^{0}$. Finally, reverting these statistically perfect changes, we obtain that $C_{i}$ is identically distributed to $\{(u_{j})^{s_{i}}(w_{j})^{X_{i,j}^{0}}\}_{j\in[m]}$, as in game $G^{*}_{5.q.3}$.
    Thus, when event $E^{\prime}$ happens, the games are identically distributed. When $E$ happens, the games both return $\beta\overset{\mathdollar}{\leftarrow}\{0,1\}$: $\mathrm{Adv}^{*}_{5.q.2}(\mathcal{B^{*}})=\mathrm{Adv}^{*}_{5.q.3}(\mathcal{B^{*}})$.
As a conclusion, the change from $G_{5.q.2}$ to $G_{5.q.3}$ is imperceptible from the point of view of the adversary $\mathcal{A}$. Therefore, $\mathrm{Adv}_{5.q.2}=\mathrm{Adv}_{5.q.3}$, for $q=1,\cdots,q_{E}$.
Theorem 2
The decentralized verifiable multi-client functional encryption for inner product supports verifiability in the random oracle, as Definition 7. More precisely, we have
$\mathrm{Adv}_{\text{CC-DVFE}}^{\text{verif}}\left(t,q_{C},q_{F}\right)\leq\epsilon_{1}+$
$q_{C}\cdot\max\left\{\mathrm{Adv}_{\prod_{Encrypt}}^{\text{snd}}(t),q_{F}\cdot\mathrm{Adv}_{\prod_{\text{DKeyGenShare}}}^{\text{snd}}(t)\right\}$
where
    $\mathrm{Adv}_{\text{CC-DVFE}}^{\text{verif}}\left(t,q_{C},q_{F}\right)$ is the best advantage of any PPT adversary running in time t against the verifiability game in Definition 7 with $q_{C}$ corruption queries and $q_{F}$ functions for the finalization phase
    $\mathrm{Adv}_{\prod_{Encrypt}}^{\text{snd}}(t)$ is the best advantage of any PPT adversary running in time $t$ against the soundness of $\prod_{\text{Encrypt}}$.
    $\mathrm{Adv}_{\prod_{\text{DKeyGenShare}}}^{\text{snd}}(t)$ is the best advantage of any PPT adversary running in time $t$ against the soundness of $\prod_{\text{DKeyGenShare}}$.
Proof 6
If we suppose that there exists a PPT adversary $\mathcal{A}$ that can win the verifiability game in Definition 7 with a non-negligible probability. As Finalize, it can be divided into the following two conditions
  • If $\mathcal{A}$ wins the game by winning the first condition with an advantage of $\epsilon_{1}$. In order to accuse an honest sender $\mathcal{C}_{i}$, $\mathcal{A}$ has to broadcast some malicious share that makes the proof of correct generation for ciphertext or functional key share of $\mathcal{C}_{i}$ invalid. By the design of the scheme, the only used elements that belongs to another client are $\left\{T_{k}\right\}_{k\in[n]}$ in the public keys, which are not used in the relation $\mathcal{R}_{Encrypt}$. As the only requirement is that $\{T_{k}\}_{k\in[n]}$ lie in the class group $\hat{G}$, submitting malformed group encodings can compromise the generation and verification processes of honest clients. However, this is a trivial attack and can be excluded, as each $\left\{T_{k}\right\}_{k\in[n],k\neq i}$ can be efficiently verified to be in the class group $\hat{G}$ in the $\prod_{\text{DKeyGenShare}}$ and the public will already know it is the corrupted client who broadcast a malicious share. Thus, the $\epsilon_{1}$ is negligible.
  • If $\mathcal{A}$ wins the game by winning the second condition. In this case, we have
    $\texttt{VerifyCT}\left((C_{\ell,i})_{i\in[n]},\pi_{CT,i},vk_{CT}\right)=1$
    and
    $\texttt{VerifyDK}\left((dk_{y,i})_{i\in[n]},\pi_{DK,i},vk_{DK},pp\right)=1.$
    This means the transcript output by $\mathcal{A}$ satisfies the relations $\mathcal{R}_{Encrypt}$ and $\mathcal{R}_{DKeyGenShare}$. From $\mathcal{R}_{\text{DKeyGenShare}}$, the tuple $(\mathrm{vk}_{\mathrm{CT}},\mathrm{vk}_{\mathrm{DK}},\mathrm{pk})$ is generated by the KeyGen algorithm, where the secret keys are given by $sk_{i}=(s_{i},\hat{k}_{i},t_{i})$, and the encryption keys are defined as $ek_{i}=s_{i}$. For each $i\in[n]$ and for each inner product function $\boldsymbol{y}$, the decryption key $dk_{y,i}$ is generated by the DKeyGenShare algorithm, taking $\left(sk_{i},\ell_{y},pk\right)$ as input. From $\mathcal{R}_{\text{Encrypt}}$, the ciphertext $C_{\ell,i}$ and the encryption proof $\pi_{\text{Encrypt},i}$ are generated by the Encrypt algorithm, using as input the message $\vec{x}_{i}\in\mathcal{M}$ and the encryption key $s^{\prime}_{i}$ for each $i\in[n]$. We model the hash function $\mathcal{H}_{1}$ as a random oracle onto $\mathbb{G}^{2}_{1}$. Then $\mathrm{com}_{i}$ is perfectly binding. From above, we have $s_{i}=s^{\prime}_{i}$. By the proved correctness of the scheme, the decryption process with input $C_{\ell}=\left\{C_{\ell,i}\right\}_{i\in[n]}$ and $h^{dk_{y}}=\text{DKeyComb}\left(\{dk_{y,i}\}_{i\in[n]},\ell_{y},pk\right)$ will output the inner product $\left\{\langle\vec{x}^{\top},y\rangle\right\}_{j\in[m]}$ for all vectors $y$, then $\mathcal{A}$ fails in the game. So $\mathcal{A}$ must either constructs a malicious ciphertext or a malicious functional key share. If $\mathcal{A}$ wins the game with a non-negligible probability, there exists a simulator $\mathcal{B}$ who can use $\mathcal{A}$ to breaking the soundness $\prod_{Encrypt}$ or $\prod_{DKeyGenShare}$. It can be constructed as follows:
      Assuming that $\mathcal{A}$ wins by constructing a malicious ciphertext with the proof. Given $q_{C}$ corrupted clients, $\mathcal{B}$ plays as a challenger in the game with $\mathcal{A}$, after $\mathcal{A}$ finalized the game, $\mathcal{B}$ guesses an index $i$ from the corrupted set $\mathcal{MC}_{\mathcal{A}}$ and outputs the instance $\left(\ell,C_{\ell,i},\mathrm{com}_{i},\pi_{i,\text{Encrypt}}\right)$ from $\mathcal{A}^{\prime}$s transcript. In the case $\mathcal{A}$ wins the game, the probability that $\mathcal{B}$ breaks the soundness of $\prod_{Encrypt}$ is $\frac{1}{q_{C}}$.
      Assuming that $\mathcal{A}$ wins by constructing a malicious functional key share with the proof. Given $q_{C}$ corrupted clients and $q_{F}$ inner-product functions $y$ to be finalized, $\mathcal{B}$ outputs one in $q_{C}\cdot q_{F}$ instances of $\left((T_{i})_{i\in[n]},d_{i},\mathrm{com}_{i},dk_{i,y}\right)$ from $\mathcal{A}$, which incurs a security loss of $q_{C}\cdot q_{F}$.
    To finalize, we have
    $\mathrm{Adv}_{\text{CC-DVFE}}^{\text{verif}}\left(t,q_{C},q_{F}\right)\leq\epsilon_{1}+$
    $q_{C}\cdot\max\left\{\mathrm{Adv}_{\prod_{Encrypt}}^{\text{snd}}(t),q_{F}\cdot\mathrm{Adv}_{\prod_{\text{DKeyGenShare}}}^{\text{snd}}(t)\right\}.$
    As $\epsilon_{1}$, $\operatorname{Adv}_{\prod_{\text{Encrypt}}}^{\text{snd}}(t)$ and $\operatorname{Adv}_{\prod_{\text{DKeyGenShare}}}^{\text{snd}}(t)$ are negligible, and $q_{C}$ and $q_{F}$ are polynomially bounded, the proof is complete.
The zero-knowledge proof algorithms related to Encrypt and VerifyCT are instantiated in Appendix B, while those related to DKeyGenShare and VerifyDK are instantiated in Appendix C. The above two zero-knowledge proofs are transformed into non-interacting zero-knowledge proofs using the Fiat-shamir transform [3], which can be applied to the CC-DVFE scheme.
VI VFEFL DESIGN
VI-A High-Level Description of VFEFL
Figure 2: The workflow and main components of VFEFL
The overview of VFEFL is depicted in the Figure 2. VFEFL utilizes the CC-DVFE to achieve a privacy and Byzantine FL among Server $\mathcal{S}$ and multiple clients $\left\{\mathcal{C}_{i}\right\}_{i\in[n]}$. Specifically, VFEFL consists of three phases. 1) Setup Phase: $\mathcal{S}$ initializes the the public parameters $pp$, the set of clients $\mathcal{CS}$, and the global initial model $W_{0}$, which are sent to $\mathcal{CS}$ (Step (1)). Each client independently generates its keys, broadcasts the corresponding the public keys and verification keys (Step (2)). 2) Model Training Phase: In the $t$-th training epoch, denoting the last global model as $W^{t-1}$, $\mathcal{S}$ updates the baseline model $W^{t}_{0}$ based on $D_{0}$ (Step (3)) and broadcasts it. Each participant downloads the $W^{t}_{0}$ and $W^{t-1}$ (Step (4)), trains a local model $W^{t}_{i}$ from $W^{t-1}$ based on the local dataset $D_{i}$ (Step (5)), encrypts it (Step (6)), and generates the functional key share (Step (7)) along with corresponding proofs, which are then sent to $\mathcal{S}$ (Step (8)). 3) Secure Aggregation Phase: $\mathcal{S}$ collects and verifies all encrypted models $\left(E_{i}\right)_{i\in\mathcal{CS}}$ and functional key shares with the proofs (Step (9)). If all verifications pass, the process moves to the next step; otherwise, the server $\mathcal{S}$ requires some clients to resend the ciphertext or key share based on the verification results and re-verifies them. $\mathcal{S}$ aggregates the authenticated functional key share, decrypts the ciphertexts, and normalizes the decryption results to get the final global update $W^{t}$ (Step (10)) and send it to the $\mathcal{CS}$.
VI-B Setup Phase
The server $\mathcal{S}$ collects all participating clients to initialize the client set $\mathcal{CS}$ and randomly sets the initial model parameters $W_{0}$. Then, the server runs the CC-DVFE.SetUp algorithm to generate the public parameters $pp$ and broadcasts $\left(\mathcal{CS},W_{0},pp\right)$. Each client $\mathcal{C}_{i}$ independently executes $C$-DVFE.KeyGen($pp$) to generate its keys pair $\left(ek_{i}=s_{i},sk_{i}=(s_{i},\hat{k}_{i},t_{i}),T_{i}\right)$, broadcasts $T_{i}$, and eventually generate the public key $pk=(T_{i},d_{i})_{i\in[n]}$ verification keys $\left(vk_{CT,i}=\mathrm{com}_{i},\ vk_{CT,i}=(T_{i},d_{i},\mathrm{com}_{i})\right)$ according to $\mathcal{CS}$. Each client sends the verification keys $vk_{CT,i}$ and $vk_{DK,i}$ to $\mathcal{S}$.
VI-C Model Training Phase
Model training consists of three parts: local training, model encryption and proof and functional key share generation and proof.
  1. 1.
    Local Training: In the $t$-th training iteration, after retrieving the global model $W^{t-1}$ from the server, each client $\mathcal{C}_{i}$ optimizes the local model using its private dataset $D_{i}$ by minimizing the corresponding loss function Loss to obtain the updated model $W_{i}^{t}$
    $W_{i}^{t}=\operatorname*{arg\min}_{W^{t}_{i}}\text{Loss}(W^{t}_{i},D_{i})$
    For benign clients, the local model is updated using stochastic gradient descent (SGD) for $R_{l}$ iterations with a local learning rate $lr$ as:
    $w^{r}\leftarrow w^{r-1}-lr\,\nabla\text{Loss}(D^{\prime}_{b};w)$
    where $r=1,..,R_{l}$, $D^{\prime}_{b}$ is a randomly sampled mini-batch from the local dataset $D_{i}$, $W_{i}^{t}=\vec{w}^{R_{l}}$. For malicious clients, they perhaps chooses a random vector as a model, or perhaps crafts it so that the global model can deviate maximally from the right direction. Similarly, $\mathcal{S}$ trains the baseline model $W^{t}_{0}$ for this epoch $t$ based on the root dataset $D_{0}$ and broadcast $W^{t}_{0}$ to the $\mathcal{C}$.
  2. 2.
[…]
  3. 3.
    Functional Key Share Generation and Proof: To ensure that malicious clients generate functional key shares honestly, we require them to provide corresponding zero-knowledge proofs during this process. For an honest client, $\mathcal{C}_{i}$ takes $y_{i}$ from the previous step and processes it using the ReLU function, resulting in $y^{\prime}_{i}=\text{ReLU}(y_{i})$. $\mathcal{C}_{i}$ executes $C$-DVFE.DKeyGenShare($sk_{i},pk,y^{\prime}_{i},\ell_{y},pp$) to generate the corresponding decryption key share $dk_{i}$ and the proof $\pi_{DK,i}$ for it. For a malicious client, $\mathcal{C}_{i}$ may try to use an invalid secret key or provide a malformed decryption key share, aiming to compromise the decryption result.
VI-D Secure Aggregation Phase
The secure aggregation phase consists of three stages: ciphertext verification , functional key share verification and aggregation.
  1. 1.
    Ciphertext Verification: In this phase, the server $\mathcal{S}$ verifies all encrypted models $\left\{E_{\ell,i}\right\}_{i\in\mathcal{C}}$ from clients using $C$-DVFE.VerifyCT($\{C_{\ell,i}\}_{i\in\mathcal{C}},\pi_{CT,i},vk_{CT}$). If all ciphertexts are valid, the process continues. Otherwise, the server identifies malicious clients $\mathcal{CC}_{CT}$, and requires all the clients in $\mathcal{CC}_{CT}$ to regenerate the ciphertext.
  2. 2.
    Functional Key Share Verification: In this phase, the server $\mathcal{S}$ collects and verifies all functional key shares from clients by executing CC-DVFE.VerifyDK ($\{dk_{y,i}\}_{i\in\mathcal{C}},\pi_{DK,i},vk_{DK},pp$). If all ciphertexts pass the verification, the process proceeds to the next phase. Otherwise, the set of malicious clients $\mathcal{CC}_{DK}$ whose key shares failed verification is identified and $\mathcal{S}$ requires all clients in $\mathcal{CC}_{DK}$ to regenerate the functional key share.
  3. 3.
    Aggregation: In this phase, the server $\mathcal{S}$ performs $\texttt{CC-DVFE.DKeyComb}\left(\{dk_{y,i}\}_{i\in\mathcal{C}},\ell_{y},pk\right)$ to aggregate the verified functional keys shares and obtain the decryption key $h^{dk_{y}}$, and $C$-DVFE.Decrypt($\left\{E_{\ell,i}\right\}_{i\in\mathcal{C}},h^{dk_{y}},y$) is executed to decrypt the verified models to gain $W^{*}$. Then $\mathcal{S}$ further normalizes $W^{*}$ to obtain $W^{t}$ according to the baseline model $W^{t}_{0}$. Finally, $\mathcal{S}$ obtains the global model $W^{t}$ for the epoch $t$ and send it to the $\mathcal{CS}$ for the next training epoch utill the global model converges.
VFEFL repeats the model training and secure aggregation phases until the global model converges, resulting in the final global model.
VII ANALYSIS
VII-A Privacy
Theorem 3
According to the security model of VFEFL, we say this VFEFL is secure if any view $\textbf{REAL}_{\mathcal{A}}^{VFEFL}$ of $\mathcal{A}$ in the real world is computationally indistinguishable from $\textbf{IDEAL}_{\mathcal{A*}}^{VFEFL}$ of $\mathcal{A}^{*}$ in the ideal world, namely
[…]
where $\hat{y}$ denotes intermediate results, $\hat{x}_{\mathcal{S}},\hat{x}_{\mathcal{C}}$ denote the inputs of $\mathcal{S}$ and $\mathcal{C}$.
Proof 7
To prove the security of VFEFL, it suffices to demonstrate the confidentiality of the privacy-preserving defense strategy, as it is the only component involving the computation of private data of clients, for the unauthorized entities $\mathcal{S}^{*}$ (the server or eavesdrops). For the curious entities $\mathcal{S}^{*}$, $\textbf{REAL}^{\mathrm{VFEFL}}_{\mathcal{A}}$ contains intermediate parameters and encrypted local models $\{E^{t}_{i}\}_{i=[n]}$ with corresponding proofs $\{\pi_{CT,i}\}_{i=[n]}$, collected during the execution of the VFEFL protocols. Additionally, we construct a PPT simulator $\mathcal{A}^{*}$ to execute $f_{\mathrm{VFEFL}}$, simulating each step of the privacy-preserving defensive aggregation strategy. The detailed proof is provided below.
Hyb1. We initialize a set of system parameters based on distributions indistinguishable from those in $\textbf{REAL}^{\mathrm{VFEFL}}_{\mathcal{A}}$ during the actual protocol execution, including public parameters $pp$, initial models $W_{0}$, and the keys.
Hyb2. In this hybrid, we alter the behavior of the simulated client $\mathcal{C}^{*}_{i}$, having it use the selected random vector $\Theta^{t}_{i}$ as its local model $W^{t}_{i}$. As only the original contents of ciphertexts have changed, the static-IND security property of VFEFL guarantees that $\mathcal{S}^{*}$ cannot distinguish the view of $\Theta^{t}_{i}$ from the view of original $W^{t}_{i}$. And in $\pi_{Encrypt,i}$, the $y_{i}$ also be sent to the server. Existing reconstruction attacks [30] cannot recover the model vector with limited information of $\|W^{t}_{i}\|$. $y_{i}$ can be viewed as a variant of $\|W^{t}_{i}\|$ and $cosine(W^{t}_{i},W^{t}_{0})$ (which is also revealed in [31]) when baseline models $W^{t}_{0}$ are known. Thus, It is a negligible security threat to $W^{t}_{i}$.
Hyb3. In this hybrid, the server uses zero-knowledge $\pi_{Encrypt}$ of encrypted random vector $\Theta^{t}_{i}$ instead of blinded models $W^{t}_{i}$ to verify the ciphertexts to obtain the result of $1$ or 0, without knowing the further information of inputs. Hence, this hybrid is indistinguishable from the previous one.
Hyb4. In this hybrid, the decryption key is computed by CC-DVFE.DKeyComb and the aggregated model $W^{*}$(before normalization) is computed by the CC-DVFE.Decrypt algorithm. $\mathcal{S}$ can obtain $\Theta^{*}=\sum_{i\in\mathcal{C}}\left(y_{i}\Theta^{t}_{i}\right)$. In addition, the local model $\Theta^{t}_{i}$ is encrypted while the randomness of these functional key shares is still preserved. This randomness ensures that the data is not compromised during the aggregation and decryption process. The probability of $\mathcal{S}$ guessing the specific $\Theta^{t}_{i}$ correctly is negligible. Hence, this hybrid is indistinguishable from the previous one.
In Hyb4, $\mathcal{A}^{*}$} holds the view $\textbf{IDEAL}_{\mathcal{A}^{*}}^{f_{\mathrm{VFEFL}}}$, which is converted from $\textbf{REAL}^{\mathrm{VFEFL}}_{\mathcal{A}}$ and is indistinguishable from it. Thus, the above analysis demonstrates that the output of $\textbf{IDEAL}_{\mathcal{A}^{*}}^{f_{\mathrm{VFEFL}}}$ is computationally indistinguishable from the output of $\textbf{REAL}_{\mathcal{A}}^{\mathrm{VFEFL}}$. Therefore, we conclude the proof of the privacy of VFEFL.
VII-B Robustness
We provide provable guarantees on the Byzantine robustness of the proposed defensive aggregation. We show that under the following assumptions, the difference between the global model $W^{t}$ learned under attack by the aggregation condition and the optimal global model $W^{opt}$ is bounded. Next, we first describe the needed assumptions and then describe our theoretical results.
Assumption 1. The expected loss function $F(w)$ is $\mu$-strongly convex and differentiable over the space $\Theta$ with an $L$-Lipschitz continuous gradient. Formally, for any $w,\hat{w}\in\Theta$, the following conditions hold:
$F(w)\geq F(\hat{w})+\langle\nabla F(\hat{w}),(w-\hat{w})\rangle+\frac{\mu}{2}\|w-\hat{w}\|^{2}.$
$\|\nabla F(w)-\nabla F(\hat{w})\|\leq L\|w-\hat{w}\|$.
Assumption 2. The local training dataset $D_{i}$ (for $i=1,2,\dots,n$) of each client and the root dataset $D_{0}$ are independently sampled from the distribution $\mathcal{X}$.
Theorem 4
For any number of malicious clients, the difference between the global model $W^{t}$ learned by VFEFL and the optimal global model $W^{opt}$ (in the absence of attacks) is bounded.
$\|W^{t}-W^{opt}\|\leq$
$3\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|W_{0}-W^{opt}\|+2\|W^{opt}\|$
where $\eta$ is the local learning rate and $W_{0}$ is the initial model. When $\|\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\|<1$, we have $\operatorname{Lim}_{t\to\infty}|W^{t}-W^{opt}\|\leq 2\|W^{opt}\|$.
Proof 8
We denote by $\mathcal{HC}$ the set of clients whose $y$ is positive in the $t$th global iteration. $W^{t}_{i}$ is the local model of client $i$ in the interaction $t$ and $W^{t}_{0}$ is the baseline model in the interaction $t$. The aggregation rule can be written as:
$W^{*}=\sum_{i\in\mathcal{HC}}\frac{\langle W^{t}_{i},W^{t}_{0}\rangle}{\langle W^{t}_{i},W^{t}_{i}\rangle}W^{t}_{i}\text{ and }W^{t}=\frac{\left\|W^{t}_{0}\right\|}{\left\|W^{*}\right\|}W^{*}$
We have the following equations for the tth global iteration:
$\|W^{t}-W^{opt}\|$
$=\left\|W^{t}-W^{t}_{0}+W^{t}_{0}-W^{opt}\right\|$
$\leq\left\|W^{t}-W^{t}_{0}\right\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$\stackrel{{\scriptstyle(a)}}{{\leq}}\left\|W^{t}+W^{t}_{0}\right\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$\leq\left\|W^{t}\right\|+\left\|W^{t}_{0}\right\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$\stackrel{{\scriptstyle(b)}}{{=}}2\left\|W^{t}_{0}\right\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$=2\left\|W^{t}_{0}-W^{opt}+W^{opt}\right\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$\leq 2\left\|W^{t}_{0}-W^{opt}\right\|+2\|W^{opt}\|+\left\|W^{t}_{0}-W^{opt}\right\|$
$=3\left\|W^{t}_{0}-W^{opt}\right\|+2\|W^{opt}\|$
$\stackrel{{\scriptstyle(c)}}{{\leq}}3\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|W_{0}-W^{opt}\|+2\|W^{opt}\|$
where $W_{0}^{t}$ is a base model trained by the server on a clean root dataset, $lr$ is the total number of server-side training iterations over $t$ rounds of iteration, (a) is because $\langle W^{t}_{i},W^{t}_{0}\rangle>0$ for $i\in\mathcal{HC}$ then $\langle W^{t},W^{t}_{0}\rangle>0$; (b) is because $\|W^{t}\|=\|W^{t}_{0}\|$; (c) is based on lemma 5. Thus, we conclude the proof.
VII-C Verifiability, self-contained and fidelity
The verifiability of VFEFL is inherently dependent on the verifiability of CC-DVFE Scheme. This dependency ensures two critical properties: it prevents malicious clients from framing honest ones, and ensures that authenticated ciphertexts and keys can be correctly decrypted. Thus, the clients successfully proves to the server that both the inputs and behavior follow the protocol specification without compromising the privacy of the model.
The proposed scheme is self-contained, as it can be deployed in the most basic federated learning framework without two non-colluding servers and any additional trusted third parties. Furthermore, it achieves high fidelity by ensuring VFEFL introduces no noise during aggregation, thereby preserving the accuracy of the global model in the absence of attacks. This fidelity is validated through experiments in Section VIII, demonstrating the scheme’s reliability in maintaining model performance.
VIII EXPERIMENTS
VIII-A Experimental Settings
In this section, we evaluate the model accuracy, Byzantine robustness and efficiency of the federated learning architecture VFEFL.
1)Implementation. We conduct our experiments on a local machine equipped with a 13th Gen Intel(R) Core(TM) i5-1340P CPU @ 1.90GHz and 16GB RAM. And we approximate floating-point numbers to two decimal places and convert them into integers for cryptographic computations, a common practice in cryptographic schemes. The final decryption step involves computing a discrete logarithm, which introduces a substantial computational overhead. Specifically, the problem reduces to solving for $x$ in the equation $a=g^{x}$, where $g$ is an element of a bilinear group and $x$ is a relatively small integer. To efficiently tackle this, we utilize the baby-step giant-step algorithm [29], which is known for its effectiveness in computing discrete logarithms in such contexts. 2)Dataset and Models. On data set allocation, we set aside a portion of the dataset as a trusted root dataset $D_{0}$ for the server and then randomly divide the remaining data into clients.
We use multiple datasets from different domains, and evaluate the accuracy of the final global model on the entire test set. The publicly available datasets are the following.
    MNIST: The dataset consists of handwritten digits, organized into 10 classes, with 60,000 samples for training and 10,000 samples for testing. Each sample is a grayscale image with a resolution of $28\times 28$ pixels.
    Fashion-MNIST: The dataset consists of grayscale images of fashion items, divided into 10 categories. It provides 60,000 training samples and 10,000 test samples, with each image having dimensions of $28\times 28$ pixels.
    CIFAR-10: This dataset consists of color images belonging to 10 distinct classes, with 50,000 training samples and 10,000 test samples. Each image has a resolution of $32\times 32$ pixels in RGB format.
3)Attack Scenario. We investigate four types of attacks: Gaussian attack, Scaling attacks, Adaptive attack and Label Flip attack. In our experiments, 20% of clients are designated as malicious under untargeted poisoning attacks, while this proportion is reduced to 10% for targeted poisoning attacks. The specific details of each attack are outlined as follows:
    Gaussian Attack[15](GA): Malicious clients may make the server’s global model less accurate by uploading random models. This attack randomly generates local models on compromised clients, which is designed using the approach and parameters from [15].
    Scaling Attack(SA): On the basis of the above Gaussian attack, the malicious client enhances the attack by amplifying its own model parameters by a scaling factor $\lambda\gg 1$ before submitting them to the server, thereby disproportionately influencing the global model. Based on prior work, we set the scaling factor $\lambda=n$, where $n$ represents the number of clients.
    Adaptive Attack[8](AA): For each aggregation rule, we define a corresponding objective function that quantifies the gap between the global model aggregated solely from honest clients and the global model after the participation of malicious clients and the solution of the optimization problem is used as a model for the attack, enabling the malicious client models to maximize the attack’s effectiveness. We use the default parameter settings in [8] and the inverse direction of the average of all models for each dimension as the initial malicious model.
    Label Flipping Attack(LFA): Our work adopts the label-flipping attack strategy outlined in [8]. Specifically, in all experiments, the malicious clients flip its label $l$ to Ml1M-l-1, where $M$ is the total number of labels and $l\in\{0,1,\cdots,M-1\}$.
4)Evaluation Metrics. In this experiment, we use the accuracy (AC) of the global model as the primary performance metric. For the LF attack, in addition to model accuracy, we also consider the attack success rate (ASR) as an evaluation criterion. The attack success rate is defined as the probability that a sample with label $l$ is misclassified as Ml1M-l-1.
VIII-B Experimental Results
VFEFL achieves the three defense goals: First, VFEFL achieves fidelity by maintaining testing error rates comparable to FedAvg under benign conditions. Second, it demonstrates robustness by effectively countering attacks from malicious clients. Finally, its efficiency remains within acceptable bounds, ensuring feasibility for real-world deployment.
First, when there is no attack, our scheme’s accuracy rates similar to FedAvg. As we can see in the Figure 3, the accuracy of our scheme is largely in line with FedAvg across multiple datasets. For instance, on MNIST, the accuracy of both FedAvg and VFEFL is close to 99%. Our aggregation rules take into account all client updates when there is no attack, so the accuracy of the global model is guaranteed. In the comparison in the table, we chose comparison schemes that are all capable of implementing fidelity protection.
(a) MNIST
(b) Fashion-MNIST
(c) CIFAR-10
Figure 3: Model accuracy on three datasets in the absence of attacks: (a) MNIST, (b) Fashion-MNIST, and (c) CIFAR-10.
Secondly, the proposed scheme demonstrates strong robustness. As illustrated in Figures 4, 5 and 6, across three different datasets, the global model trained under our scheme consistently maintains high accuracy when subjected to Gaussian attack, Scaling attack, and Adaptive attack. In the case of Gaussian attacks, the accuracy of the global model drops by less than 0.1%. The other three baseline methods also exhibit strong robustness against Gaussian attack, however, they fail to defend effectively against Scaling attacks. This vulnerability arises from the fact that these methods neither constrain nor validate the structural characteristics of local models submitted by malicious clients. As a result, when adversaries employ models with sufficiently large magnitudes, they can exert a high influence on the final aggregated model. In contrast, thanks to the robust aggregation rule in our scheme, which effectively verifies and restricts the norm of local models continues to sustain high accuracy even in the presence of Scaling attacks. In the presence of a Adaptive attack, both BSR-FL and VFEFL are successfully aggregated into a final global model that achieves sufficiently high accuracy. Overall, these experimental results validate that our federated learning framework possesses strong robustness against various types of untargeted poisoning attacks.
(a) GA
(b) SA
(c) AA
Figure 4: Model accuracy under different attacks on MNIST: (a) GA, (b) SA, and (c) AA.
(a) GA
(b) SA
(c) AA
Figure 5: Model accuracy under different attacks on Fashion-MNIST: (a) GA, (b) SA, and (c) AA.
(a) GA
(b) SA
(c) AA
Figure 6: Model accuracy under different attacks on CIFAR-10: (a) GA, (b) SA, and (c) AA.
As illustrated in Figures 7, 8 and 9, the proposed scheme consistently achieves high accuracy under LF attacks across all three datasets, while effectively suppressing the attack success rate. This highlights the robustness of our approach in defending against LF attacks. But in the CIFAR10 dataset, the global models obtained from the training of our scheme converge relatively slowly, as shown in Fig 9. This phenomenon is mainly due to the fact that the adopted aggregation rule normalises the norm of each local model based on the root model. This design effectively suppresses the malicious client’s strategy to enhance the attack effectiveness by amplifying the parameters of the local models, as demonstrated by the scaling attack in Figs 4b, 5b, and 6b. However, this normalisation mechanism also increases the number of iterations required for model convergence to some extent. Nevertheless, after about 20 iterations, the global model trained by our scheme is still able to achieve satisfactory accuracy. In addition, the global model trained by the proposed scheme not only maintains high classification accuracy on MNIST and Fashion-MNIST datasets, but also demonstrates better defence capability and performance stability against malicious attacks compared to other schemes in multiple iterations. The above results further validate the robustness and effectiveness of this paper’s method in malicious environments.
(a) AC under LF attacks
(b) ASR under LF attacks
Figure 7: AC and ASR under LF attacks on MNIST: (a) accuracy (AC) and (b) attack success rate (ASR).
(a) AC under LF attacks
(b) ASR under LF attacks
Figure 8: AC and ASR under LF attacks on Fashion-MNIST: (a) accuracy (AC) and (b) attack success rate (ASR).
(a) AC under LF attacks
(b) ASR under LF attacks
Figure 9: AC and ASR under LF attacks on CIFAR-10: (a) accuracy (AC) and (b) attack success rate (ASR).
Finally, we conduct a comprehensive evaluation of the computational efficiency of the federated learning framework. First, a performance test of the constructed CC-DVFE scheme was conducted to measure the time consumed for each encryption and decryption as well as the generation and verification of ciphertexts and keys, and the results are shown in Table III. The time required for the Setup phase exhibits some uncertainty due to the need for trial and error in generating appropriate prime numbers within the GenClassGroup. The runtime of KeyGen abd DKeyGenShare scales linearly with the number of clients, while that of Encryption and VerifyCT increases with the ciphertext dimensionality. In the Encryption, hash to point operation is relatively time-consuming, however, it only needs to be executed once during the system construction. In this experiment, the encryption, key generation, and related verification operations are recorded as the average time for each client.
TABLE III: Time Consumption of CC-DVFE Under Different Dimensions
Operations m = 10 m = 50 m = 100 m = 500 m = 1000
Setup 9.5959 s 4.8450 s 9.3008 s 14.8908 s 10.5898 s
Key Generation 1.3672 s 1.3086 s 1.3185 s 1.3033 s 1.2764 s
Encryption Hash to point 0.0242 s 0.1144 s 0.2174 s 1.0573 s 2.0880 s
Encryption 0.0077 s 0.0189 s 0.0427 s 0.2123 s 0.4761 s
Ciphertext Proof 0.0235 s 0.0628 s 0.1080 s 0.4817 s 1.0968 s
Verify Ciphertext 0.0346 s 0.0749 s 0.1240 s 0.4899 s 1.1786 s
DKeyGenShare Key Generation 0.0186 s 0.0189 s 0.0184 s 0.0181 s 0.0197 s
Key Proof 0.1401 s 0.1523 s 0.1378 s 0.1351 s 0.1477 s
Verify Functional Key 0.2237 s 0.2276 s 0.2174 s 0.2163 s 0.2361 s
Key Combination 0.0165 s 0.0204 s 0.0158 s 0.0158 s 0.0157 s
Decryption 0.1400 s 0.9296 s 1.4903 s 7.0142 s 13.2858 s
TABLE IV: Time Consumption of VFEFL Per Iteration (LeNet5, MNIST, $n=10$, $m=61706$)
Operations Time (s)
Setup Phase Setup 87.54
Communication 0.43
Model Training Phase Local Training 99.39
Hash to Pairing 128.61
Model Encryption and Proof 87.55
Key Share Generation and Proof 0.16
Communication 12.38
Secure Aggregation Phase Ciphertext Verification 593.93
Key Share Verification 2.35
Aggregation 1014.62
Based on the results of this test, we further evaluate the time required to complete a full training iteration using the MNIST dataset under the VFEFL framework, and the relevant results are demonstrated in Table IV. In practical experiments, the Setup phase also includes some preprocessing operations related to solving the discrete logarithm problem. The communication time of the Setup Phase refers to the time taken for the server to send the initial model and the baseline model to the client. The communication time of the Model Training Phase refers to the time taken for the client to upload the encrypted model, the decryption key share, and corresponding proofs to the server. These communication times are averaged over multiple clients. In the aggregation phase, the server performs the verification of all ciphertexts and keys, and the total verification time is computed. On the whole, despite the introduction of cryptographic operations, the overall training efficiency of the federated learning framework is still within an acceptable range, which verifies that the proposed scheme can effectively take into account the computational performance while guaranteeing security.
IX CONCLUSION
In this paper, we presented VFEFL, based on decentralized verifiable functional encryption. To implement the framework, we proposed a new aggregation rule and a new verifiable functional encryption scheme, and proved their security. Our VFEFL scheme achieves privacy protection and resistance to Byzantine clients in a single-server setup without introducing additional trusted third parties, and is more applicable to general scenarios. We analyzed both the privacy and robustness of the scheme, in addition to the verifiability, fidelity and self-containment achieved by VFEFL. Finally, we further demonstrated the effectiveness of the VFEFL scheme through experiments.
Appendix A assumptions and security models
Definition 2 (The Symmetric eXternal Diffie-Hellman Assumption)
SXDH Assumption states that, in a pairing group $\mathcal{PG}$, the DDH assumption holds in both $\mathbb{G}_{1}$ and $\mathbb{G}_{2}$.
Definition 3 (HSM assumption)
Let $\text{GenClassGroup}=(\texttt{Gen},\texttt{Solve})$ be a generator for DDH groups with an easy DL subgroup. Let $(\tilde{s},f,\hat{h}_{p},\hat{G},F)$ be the output of $\text{Gen}(\lambda,p)$, and set $h:=h_{p}f$. Denote by $D$ (resp. $\mathcal{D}_{p}$) a distribution over integers such that $\{h^{x},x\leftarrow D\}$ (resp. $\{\hat{h}_{p}^{x},x\leftarrow\mathcal{D}_{p}\}$) is within statistical distance $2^{-\lambda}$ from the uniform distribution in $\langle h\rangle$ (resp. $\langle\hat{h}_{p}\rangle$). For any adversary $\mathcal{A}$ against the HSM problem, its advantage is defined as:
$\text{Adv}_{\mathcal{A}}^{\text{HSM}}(\lambda):=$
$\left|\Pr\left[\begin{array}[]{l}\begin{matrix}(\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p})\\ \leftarrow\texttt{Gen}(\lambda,p),\\ t\leftarrow\mathcal{D}_{p},\,h_{p}=\hat{h}_{p}^{t},\,x,x^{\prime}\\ \leftarrow D,\,b\overset{\mathdollar}{\leftarrow}\{0,1\},\\ Z_{0}=h^{x},\,Z_{1}=h^{x^{\prime}},\,b^{\prime}\leftarrow\\ \mathcal{A}(p,\tilde{s},f,\hat{h}_{p},h_{p},F,\hat{G}^{p},Z_{b},\texttt{Solve}(\cdot))\end{matrix}:b=b^{\prime}\end{array}\right]-\frac{1}{2}\right|.$
The HSM problem is hard in $G$ if for all probabilistic polynomial-time adversaries $\mathcal{A}$, $\text{Adv}_{\mathcal{A}}^{\text{HSM}}(\lambda)$ is negligible.
From [25], one can set $S=2^{\lambda-2}\tilde{s}$, and instantiate $\mathcal{D}_{p}$ as the uniform distribution on $\{0,\cdots,S\}$ and $\mathcal{D}$ as the uniform distribution on $\{0,...,pS\}$.
Definition 4 (Decisional Diffie-Hellman Assumption)
The Decisional Diffie-Hellman (DDH) assumption states that, in a prime-order group $\mathcal{G}\leftarrow\text{GGen}(\lambda)$, no probabilistic polynomial-time (PPT) adversary can distinguish between the following two distributions with non-negligible advantage, namely
$\mathrm{Adv}_{\mathbb{G}}^{ddh}(t)=|$ $\Pr\left[(g^{a},g^{b},g^{ab})\mid a,b\leftarrow\mathbb{Z}_{p}\right]$
$-\Pr\left[(g^{a},g^{b},g^{c})\mid a,b,c\leftarrow\mathbb{Z}_{p}\right]|<\epsilon(\lambda)$
Definition 5 (Multi DDH Assumption[14])
For any distinguisher $\mathcal{A}$ with runtime at most $t$, its maximum advantage in distinguishing the following two distributions:
$\mathcal{D}_{m}=\left\{\left(g^{x},(g^{y_{j}},g^{xy_{j}})_{j\in[m]}\right)\mid x,y_{j}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\right\},$
$\mathcal{D}^{\prime}_{m}=\left\{\left(g^{x},(g^{y_{j}},g^{z_{j}})_{j\in[m]}\right)\mid z_{j}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}\right\}$
is upper bounded by
$\operatorname{Adv}^{\mathrm{multi-ddh}}_{\mathbb{G}}(t)\leq\operatorname{Adv}^{\mathrm{ddh}}_{\mathbb{G}}(t+4m\cdot t_{\mathbb{G}}),$
where $t_{\mathbb{G}}$ denotes the time for an exponentiation in $\mathbb{G}$.
Definition 6 (Static-IND-Security for CC-DVFE)
Consider a CC-DVFE scheme applied to a set of $n$ clients, each associated with an $m$-dimensional vector. No PPT adversary $\mathcal{A}$ should be able to win the following security game with at least two non-corrupted clients, against a challenger $\mathcal{C}$ with non-negligible probability.
  • Initialization: The challenger $\mathcal{C}$ runs algorithm SetUp($\lambda$) and KeyGen($pp$) to get public parameters $pp$ and the keys $\left((sk_{i},ek_{i})_{i\in[n]},vk_{CT},vk_{DK},pk\right)$ and chooses a random bit $b\overset{\mathdollar}{\leftarrow}\{0,1\}$. It sends $\left(vk_{CT},vk_{DK},pk\right)$ to the $\mathcal{A}$. $\mathcal{A}$ submits the indexs $\mathcal{CC}$ of the corruption queries for QCorrupt.
[…]
  • Corruption queries QCorrupt: $\mathcal{A}$ can corruption queries by submitting the index $id\in\mathcal{CC}$ up to $q_{C}$ times. $\mathcal{C}$ returns secret and encryption keys $(sk_{id},ek_{id})$ to $\mathcal{A}$.
  • Finalize: $\mathcal{A}$ outputs its guess $b^{\prime}$ on the bit $b$, and this process produces the outcome $\beta$ of the security game, based on the analysis provided below. Let $\mathcal{CC}$ represent the set of corrupted clients (the indices $id$ that are input to QCorrupt throughout the entire game), while $\mathcal{HC}$ denotes the set of honest (noncorrupted) clients. We set output $\beta=(b^{\prime}==b)$ if none of the following occur, otherwise, $\beta\overset{\mathdollar}{\leftarrow}\{0,1\}$
    1. 1.
      A QEncrypt($id,X^{0}_{id},X^{1}_{id},\ell,\left\{\ell_{Enc}\right\}_{j\in[m]}$)-query has been made for an index $(id\in\mathcal{CC})$ where $(X^{0}_{i}\neq X^{1}_{i})$.
    2. 2.
[…]
    3. 3.
      For a given label $\ell$ and a function vector $\vec{y}$ asked for QDKeyGen, for all $id\in\mathcal{CC}$, $X_{id}^{0}=X_{id}^{1}$ and for all $id^{\prime}\in\mathcal{HC}$, QEncrypt($id^{\prime},X^{0}_{id^{\prime}},X^{1}_{id^{\prime}},\ell,\left\{\ell_{Enc}\right\}_{j\in[m]}$) have been made, but there exists $j\in[m]$ such that $\langle X^{\top 0}_{j},\vec{y}\rangle\neq\langle X^{\top 1}_{j},\vec{y}\rangle$ holds, where $X^{\top 0}_{j}=\{X^{0}_{1,j},\cdots,X^{0}_{n,j}\}\in\mathcal{M}^{n}$, $X^{\top 1}_{j}=\{X^{1}_{1,j},\cdots,X^{1}_{n,j}\}\in\mathcal{M}^{n}$
    4. 4.
[…]
We say the CC-DVFE is $\epsilon(\lambda)$-static-IND-secure if there exists no adversary $\mathcal{A}$ who can win the above game with advantage at least $\epsilon(\lambda)$. Namely,
$\left|\mathrm{Adv_{CC-DVFE}^{sta-ind}}\right|=\left|\mathrm{Pr}[b^{\prime}=b]-\frac{1}{2}\right|\leq\epsilon(\lambda)$
Definition 7 (Verifiability for CC-DVFE)
Consider a CC-DVFE scheme applied to a set of $n$ clients, each associated with an $m$-dimensional vector. No PPT adversary $\mathcal{A}$ should be able to win the following security game against a challenger $\mathcal{C}$ with non-negligible probability.
  • Initialization: Challenger $\mathcal{C}$ initializes by running $\texttt{SetUp}(\lambda)\to pp$. It sends $pp$ to $\mathcal{A}$.
  • Key generation queries QKeyGen: For only one time in the game, $\mathcal{A}$ can play on behalf of corrupted clients $\mathcal{CC}$ to join a KeyGen($pp$). The challenger $\mathcal{C}$ executes KeyGen($pp$) on representations of non-corrupted clients $\mathcal{HC}$ and finally output $(vk_{CT},vk_{DK},pk)$.
  • Corruption queries QCorrupt: $\mathcal{A}$ can possesses $q_{C}$ access to the corruption queries with submitting an index $id$ to play on behalf of client $id$ in the protocol. $\mathcal{C}$ sends $(sk_{id},ek_{id},vk_{CT},vk_{DK})$ to $\mathcal{A}$ if $id$ is queried after QKeyGen. And $\mathcal{A}$ cannot play on behalf of client $id$ in the key generation process anymore.
[…]
  • Finalize: Let $\mathcal{CC}$ be the set of corrupted clients and $\mathcal{HC}$ denotes the set of honest clients, then $\mathcal{A}$ has to output verification keys $(vk_{CT},vk_{DK})$ and public key $pk$ from QKeyGen(), a set of labels $\ell$ and $\left(\ell_{Enc,j}\right)_{j\in[m]}$ , malicious ciphertexts $(C_{\ell,id},\pi_{CT,id})_{id\in\mathcal{CC}}$, and malicious functional key shares $(dk_{y,id},\pi_{DK,id})_{id\in\mathcal{CC}}$ for a polynomially number $q_{F}$ of function vectors $\{y_{f}\}_{f\in[q_{F}]}$. The ciphertexts of honest clients $\{C_{\ell,id}\}_{id\in\mathcal{HC}}$ and their functional key shares $\{dk_{y,id}\}_{id\in\mathcal{HC}}$ are automatically completed by using the oracles QEncrypt and QDKeyGen. $\mathcal{A}$ wins the game if one of the following cases happens:
    1. 1.
      If $\texttt{VerifyCT}((C_{\ell,id})_{id\in[n]},\pi_{CT,id},vk_{CT})\to 0$ or $\texttt{VerifyDK}((dk_{id,y})_{id\in[n]},\pi_{DK,i},vk_{DK},pp)\to 0$, for some $y_{f}$, the union $\mathcal{CC}=\mathcal{CC}_{CT}\cup\mathcal{CC}_{DK}$ contains an honest sender $\mathcal{C}_{i}$. In other words, an honest client is falsely accused of being a corrupt client, which is noted as event E1.
    2. 2.
      If $\texttt{VerifyCT}((C_{\ell,id})_{id\in[n]},\pi_{\text{CT},i},vk_{\text{CT}})\to 1$ and $\texttt{VerifyDK}((dk_{id,y})_{id\in[n]},\pi_{\text{DK},i},vk_{\text{DK}},pp)\to 1$ there does not exist a tuple of messages $(\vec{x}_{id})_{id\in[n]}$ such that $\texttt{Decrypt}((C_{\ell,id})_{id\in[n]},{dk},\vec{y})\to(\langle\vec{x}^{\top},\vec{y}\rangle)_{j\in[m]}$ holds where $\vec{x}^{\top}=(x_{1,j},\dots,x_{n,j})$$\in\mathcal{M}^{n}$ for all function vectors $y_{f}$ with $f\in[q_{F}]$, where $dk=\texttt{DKeyComb}(\{dk_{i,y_{f}}\}_{i\in[n]},\ell_{y},pk)$. This defines event E2.
    The first winning case guarantees that no honest client will be misjudged when authentication fails. The second winning condition is determined statistically, and the scheme must ensure that the adversary cannot generate a ciphertext and decryption key that can be verified, causing the decryption to fail.
We say the CC-DVFE has $\epsilon(\lambda)$-verifiability if there exists no adversary $\mathcal{A}$ who can win the above game with advantage at least $\epsilon(\lambda)$. Namely,
$\left|\mathrm{Adv_{CC-DVFE}^{ver}}\right|=\left|\text{Pr}(\text{E1})+\text{Pr}(\text{E2})\right|\leq\epsilon(\lambda).$
Appendix B Instantiation of $\pi_{CT}$
In the two algorithms Encrypt and VerifyCT of the above CC-DVFE scheme, we implement the verification through the following zero-knowledge proof. For label $\ell$ and encryption label $\ell_{Enc}$, the vector $\vec{x}_{i}\in\mathcal{M}^{m}$, auxiliary vector $\vec{x}_{0}\in\mathcal{M}^{m}$ and commit of key $\mathrm{com}_{i}$, we define the relation:
$\mathcal{R}_{Encrypt}(V_{i},\mathrm{com}_{i},\ell,\ell_{Enc},y_{i},\vec{x}_{0})=1\leftrightarrow\begin{cases}V_{i}=(\bar{u})^{s_{i}}\cdot\vec{w}^{\vec{x}_{i}}\\ y_{i}=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}\\ \mathrm{com}_{i}=(v)^{s_{i}}\end{cases}.$
where $\bar{u}=\prod_{j}^{m}u_{j}$, $u_{j}=\mathcal{H}_{1}(\ell||j)$, $\vec{w}^{{x}_{i}}=\prod_{j\in[m]}(w_{j})^{x_{i,j}}$, $\{w_{j}\}_{j\in[m]}=\{\mathcal{H}_{1}^{{}^{\prime}}(\ell_{Enc,j}||j)\}_{j\in[m]}$, $v=\mathcal{H}_{1}(\ell_{D})$.
We proposes a zero knowledge proof system for the relation as described above:
$\text{PoK}\left\{(s_{i},\vec{x}_{i}):\begin{aligned} V_{i}=(\bar{u})^{s_{i}}\cdot\vec{w}^{\vec{x}_{i}}\wedge\\ y_{i}=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}.\end{aligned}\right\}$
To prove the statement, the prover $\mathcal{P}$ and $\mathcal{V}$ engage in the following zero knowledge protocol.
$\mathcal{P}$ computes:
$\rho\overset{\mathdollar}{\leftarrow}\mathbb{Z}^{2}_{p},\quad\vec{k}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{m}$
$K=(\bar{u})^{\rho}\vec{w}^{\vec{k}},\quad L(X)=\vec{x}_{i}+\vec{k}X$
$t_{0}=\langle\vec{x}_{i},\vec{x}_{i}\rangle,\quad t_{1}=2\langle\vec{k},\vec{x}_{i}\rangle,\quad t_{2}=\langle\vec{k},\vec{k}\rangle$
$t(X)=\langle L(X),L(X)\rangle=t_{0}+t_{1}X+t_{2}X^{2}$
$\tau_{0},\ \tau_{1},\ \tau_{2},\ \tau^{\prime}_{1}\leftarrow\mathbb{Z}^{2}_{p}$
$T_{0}=g^{t_{0}}(\bar{u})^{\tau_{0}},\quad T_{1}=g^{t_{1}}(\bar{u})^{\tau_{1}},\quad T_{2}=g^{t_{2}}(\bar{u})^{\tau_{2}}$
$t^{\prime}_{1}=\langle\vec{k},\vec{x}_{0}\rangle,\quad T^{\prime}_{1}=g^{t^{\prime}_{1}}(\bar{u})^{\tau^{\prime}_{1}}$
$\mathrm{com}^{*}=(v)^{\rho},\quad v=\mathcal{H}_{1}(\ell_{D})$
$\mathcal{P}\to\mathcal{V}:K,T_{0},T_{1},T_{2},T^{\prime}_{1},\mathrm{com}^{*}_{i}$
$\mathcal{V}:\alpha\overset{\mathdollar}{\leftarrow}\mathbb{Z}^{*}_{p}$
$\mathcal{V}\to\mathcal{P}:\alpha$
$\mathcal{P}$ computes:
$\tau_{\alpha}=\tau_{0}+\tau_{1}\alpha+\tau_{2}\alpha^{2},\quad\tau^{\prime}_{\alpha}=\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha$
$\omega=s_{i}+\alpha\rho\in\mathbb{Z}^{2}_{p},\quad L=L(\alpha),\quad t=t(\alpha)$
$\mathcal{P}\to\mathcal{V}:\tau_{\alpha},\tau^{\prime}_{\alpha},\omega,L,t$
$\mathcal{V}$ verifies:
$g^{t}(\bar{u})^{\tau_{\alpha}}$ $\overset{?}{=}T_{0}\left(T_{1}\right)^{\alpha}\left(T_{2}\right)^{\alpha^{2}}$
$g^{\langle L,\vec{x}_{0}\rangle}(\bar{u})^{\tau^{\prime}_{\alpha}}$ $\overset{?}{=}T_{0}^{y_{i}}(T^{\prime}_{1})^{\alpha}$
$\vec{w}^{L}(\bar{u})^{\omega}$ $\overset{?}{=}V_{i}(K)^{\alpha}$
$t$ $\overset{?}{=}\langle L,L\rangle$
$(v)^{\omega}$ $\overset{?}{=}\mathrm{com}_{i}(\mathrm{com}^{*})^{\alpha}$
Theorem 5
The Ciphertext Verification $\prod_{Encrypt}$ has perfect completeness, perfect honest verifier zero-knowledge and computational special soundness.
Proof 9
A formal proof are provided in Appendix D.
Appendix C Instantiation of $\pi_{DK}$
In the two algorithms DKeyGenShare and VerifyDK of the above CC-DVFE scheme, we implement the verification through the following zero-knowledge proof. To prevent malicious clients from interfering with the aggregation process by uploading malicious keys, we define the relation:
$\mathcal{R}_{DKeyGenShare}(\left(T_{i}\right)_{i\in[n]},d_{i},dk_{i},\ell_{y},y_{i})=1$
$\leftrightarrow\begin{cases}T_{i}=\left(h_{p}^{t_{i,b}}\right)_{b\in[2]}\\ d_{i}=\left(f^{\hat{k}_{b}}K_{\sum,i,b}^{t_{i,b}}\right)_{b\in[2]}\\ dk_{i}=\left((\hat{v}_{b})^{\hat{k}}\cdot h^{s_{i,b}\cdot y_{i}}\right)_{b\in[2]}\\ \mathrm{com}_{i}=(v)^{s_{i}}\end{cases}.$
where $K_{\sum,i,b}=\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\in G$, $\hat{v}_{b}=\mathcal{H}_{2}(\ell_{y,b}),b\in[2],v=\mathcal{H}_{1}(\ell_{D})$.
We guarantee the correctness of the decryption key share by constructing zero-knowledge proofs as:
$\text{PoK}\left\{(s_{i},t_{i},\hat{k}):\begin{aligned} T_{i}=h_{p}^{t_{i}}\wedge\\ d_{i}=(f^{\hat{k}_{b}}K_{\sum,i,b}^{t_{i,b}})_{b\in[2]}\wedge\\ dk_{i}=\left((\hat{v}_{b})^{\hat{k}}\cdot h^{s_{i,b}\cdot y_{i}}\right)_{b\in[2]}\wedge\\ \mathrm{com}_{i}=(v)^{s_{i}}\end{aligned}\right\}$
To prove the statement, the prover $\mathcal{P}$ and $\mathcal{V}$ engage in the following zero knowledge protocol.
$\mathcal{V}$ verifes that $d_{i},T_{i}\in\hat{G}^{2}$ for $j\in[n]$.
$\mathcal{P}$ computes:
$r_{\hat{k},i},r_{s,i}\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}^{2},\quad r_{t,i}\overset{\mathdollar}{\leftarrow}[0,2^{\lambda}pS]^{2}$
$R_{T,i}=(h^{r_{t,i,b}}_{p})_{b\in[2]}$
$K_{\sum,i,b}=\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\in G$
$R_{d,i}=\left(f^{r_{\hat{k},i,b}}K_{\sum,i,b}^{r_{t,i,b}}\right)_{b\in[2]}$
$R_{dk,i}=\left((\hat{v}_{b})^{r_{\hat{k},i}}\cdot h^{r_{s,i.b}\cdot y_{i}}\right)_{b\in[2]}$
$R_{com,i}=(v)^{r_{s,i}}$
$\mathcal{P}\to\mathcal{V}:R_{T,i},R_{d,i},R_{dk,i},R_{com,i}$
$\mathcal{V}:\beta\overset{\mathdollar}{\leftarrow}\mathbb{Z}_{p}$
$\mathcal{V}\to\mathcal{P}:\beta$
$\mathcal{P}$ computes:
$z_{\hat{k},i}=r_{\hat{k},i}-\beta\cdot\hat{k}_{i}$
$z_{s,i}=r_{s,i}-\beta\cdot s_{i}$
$z_{t,i}=r_{t,i}-\beta\cdot t_{i}$
$\mathcal{P}\to\mathcal{V}:z_{\hat{k},i},z_{s,i},z_{t,i}$
$\mathcal{V}$ verifies:
$R_{T,i}$ $\overset{?}{=}(T_{i,b}^{\beta}\cdot h_{p}^{z_{t,i,b}})_{b\in[2]}$
$K_{\Sigma,i}$ ${=}\left\{\prod_{i<i^{*}}^{n}T_{i^{*},b}\cdot\left(\prod_{i>i^{*}}^{n}T_{i^{*},b}\right)^{-1}\right\}_{b\in[2]}\in G$
$R_{d,i}$ $\overset{?}{=}\left(\mathrm{d}_{i,b}^{\beta}\cdot f^{z_{\hat{k},i,b}}K_{\Sigma,i,b}^{z_{t,i,b}}\right)_{b\in[2]}$
$R_{\mathrm{dk}_{i}}$ $\overset{?}{=}\left((\mathrm{dk}_{i})^{\beta}\cdot(\hat{v}_{b})^{z_{\hat{k},i}}\cdot h^{z_{s,b}\cdot y_{i}}\right)_{b\in[2]}$
$R_{\mathrm{com},i}$ $\overset{?}{=}(\mathrm{com}_{i})^{\beta}(v)^{z_{s,i}}$
Theorem 6
The functional key share verification $\prod_{DKeyGenShare}$ has perfect completeness, statistically zeroknowledge, and a computational soundness.
Proof 10
A formal proof are provided in [25].
Appendix D Proof of Theorem 5
Proof 11
Completeness. Perfect completeness follows from the fact that for all valid witnesses, it holds that $y_{i}=\frac{\langle\vec{x}_{i},\vec{x}_{0}\rangle}{\langle\vec{x}_{i},\vec{x}_{i}\rangle}$ and
$g^{t}(\bar{u})^{\tau_{\alpha}}$ $=g^{t_{0}+t_{1}\alpha+t_{2}\alpha^{2}}(u)^{\tau_{0}+\tau_{1}\alpha+\tau_{2}\alpha^{2}}$
$=g^{t_{0}}u^{\tau_{0}}(g^{t_{1}}u^{\tau_{1}})^{\alpha}(g^{t_{2}}u^{\tau_{2}})^{\alpha^{2}}$
$=T_{0}\left(T_{1}\right)^{\alpha}\left(T_{2}\right)^{\alpha^{2}}$
$g^{\langle L,\vec{x}_{0}\rangle}(\bar{u})^{\tau^{\prime}_{\alpha}}$ $=g^{\langle\vec{x}_{i}+k\alpha,\vec{x}_{0}\rangle}u^{\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha}$
$=(g^{t_{0}}u^{\tau_{0}})^{y_{i}}(g^{t^{\prime}_{1}}(u)^{\tau^{\prime}_{1}})^{\alpha}$
$=T_{0}^{y_{i}}(T^{\prime}_{1})^{\alpha}$
$\vec{w}^{L}(\bar{u})^{\omega}$ $=\vec{w}^{\vec{x}_{i}+k\alpha}u^{s_{i}+\alpha\rho}$
$=\vec{w}^{\vec{x}_{i}}u^{s_{i}}(\vec{w}^{k}u^{\rho})^{\alpha}=V_{i}(K)^{\alpha}$
$t$ $=\langle L,L\rangle$
$(v)^{\omega}$ $=v^{s_{i}+\alpha\rho}=v^{s_{i}}(v^{\rho})^{\alpha}=\mathrm{com}_{i}(\mathrm{com}^{*})^{\alpha}.$
Soundness. To demonstrate special soundness, we construct an extractor $\mathcal{E}$ as outlined below. $\mathcal{E}$ rewinds the $\prod_{\texttt{Encrypt}}$ with $2$ different values of $\alpha$ to get
$L_{1}=\vec{x}_{i}+k\alpha_{1}$ $\text{ and }L_{2}=\vec{x}_{i}+k\alpha_{2}$
$\omega_{1}=s_{i}+\alpha_{1}\rho$ $\text{ and }\omega_{2}=s_{i}+\alpha_{2}\rho$
$\tau^{\prime}_{\alpha_{1}}=\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha_{1}$ $\text{ and }\tau^{\prime}_{\alpha_{2}}=\tau_{0}y_{i}+\tau^{\prime}_{1}\alpha_{2}$
$\tau_{\alpha_{1}}=\tau_{0}+\tau_{1}\alpha_{1}+\tau_{2}\alpha_{1}^{2}$ $\text{ and }\tau_{\alpha_{2}}=\tau_{0}+\tau_{1}\alpha_{2}+\tau_{2}\alpha_{2}^{2},$
then $\mathcal{E}$ can compute
$k$ $=\frac{L_{1}-L_{2}}{\alpha_{1}-\alpha_{2}},\quad\rho=\frac{\omega_{1}-\omega_{2}}{\alpha_{1}-\alpha_{2}}$
$\vec{x}_{i}$ $=L_{1}-k\alpha_{1}=L_{1}-\alpha_{1}\frac{L_{1}-L_{2}}{\alpha_{1}-\alpha_{2}}$
$s_{i}$ $=\omega_{1}-\alpha_{1}\rho=\omega_{1}-\alpha_{1}\frac{\omega_{1}-\omega_{2}}{\alpha_{1}-\alpha_{2}}$
$\tau^{\prime}_{1}$ $=\frac{\tau^{\prime}_{\alpha_{1}}-\tau^{\prime}_{\alpha_{2}}}{\alpha_{1}-\alpha_{2}},\quad\tau_{0}=\frac{\tau^{\prime}_{\alpha_{1}}-\tau^{\prime}_{1}\alpha_{1}}{y_{i}}$
$\tau_{1}$ $=\frac{\alpha_{2}(\tau^{\prime}_{\alpha_{1}}-\tau_{0})}{\alpha_{1}(\alpha_{2}-\alpha_{1})}-\frac{\alpha_{1}(\tau_{\alpha_{2}}-\tau_{0})}{\alpha_{2}(\alpha_{2}-\alpha_{1})}$
$\tau_{2}$ $=\frac{\tau_{\alpha_{1}}-\tau_{0}}{\alpha_{1}(\alpha_{1}-\alpha_{2})}-\frac{\tau_{\alpha_{2}}-\tau_{0}}{\alpha_{2}(\alpha_{1}-\alpha_{2})}$
to extract $\vec{x}_{i},s_{i},\tau_{0},\tau_{1},\tau_{2},\tau^{\prime}_{1}$, and then further computation of
$t_{0}=\langle\vec{x}_{i},\vec{x}_{i}\rangle,\ t_{1}=2\langle k,\vec{x}_{i}\rangle,\ t_{2}=\langle k,k\rangle,\ t^{\prime}_{1}=\langle k,\vec{x}_{0}\rangle$
reveals that $t_{0},t_{1},t_{2},t_{1}^{\prime}$.
Zero knowledge. To demonstrate the protocol achieves perfect honest-verifier zero-knowledge, we construct a simulator that, given a statement $\left(V_{i}\in\mathbb{G}_{1},y_{i}\in\mathbb{Z}_{p},\mathrm{com}_{i}\in\mathbb{G}^{2}_{1}\right)$, produces a simulated proof whose distribution is indistinguishable from that of a real proof generated through interaction between an honest prover and an honest verifier. All elements chosen by the simulator are either independently and uniformly distributed or fully determined by the verification equations. The simulator randomly chooses $T_{0},T_{2},\tau_{\alpha},\tau^{\prime}_{\alpha},\omega$ and $L$. Then, for the challenge $\alpha$ and the public value $\vec{x}_{0}$, the values $t,T_{1},T^{\prime}_{1},K$, and $\mathrm{com}^{*}$ are computed according to the verification equations.
$t$ $=\langle L,L\rangle$
$T_{1}$ $=(g^{t}(\bar{u})^{\tau_{\alpha}}T_{0}^{-1}T_{2}^{-\alpha^{2}})^{-\alpha}$
$T^{\prime}_{1}$ $=\left(g^{\langle L,\vec{x}_{0}\rangle}(\bar{u})^{\tau^{\prime}_{\alpha}}T_{0}^{-y_{i}}\right)^{-\alpha}$
$K$ $=\left(\vec{w}^{L}(\bar{u})^{\omega}(V_{i})^{-1}\right)^{-\alpha}$
$\mathrm{com}^{*}$ $=(v^{\omega}(\mathrm{com}_{i})^{-1})^{-\alpha}$
The protocol remains zero-knowledge because we can successfully simulate the proof procedure. Therefore, any information revealed or potentially leaked during the simulation does not compromise the zero-knowledge property of the overall protocol.
Appendix E Lemma 5 and Proof
Lemma 5
Assume the objective function $F(W)$ is $\mu$-strongly convex and $L$-smooth (Assumption 1). Let $W^{opt}$ denote the optimal global model of $F(W;D)$, and $W^{t}_{0}$ denote the baseline model at iteration $t$, obtained using gradient descent on a clean dataset $D_{0}$ (Assumption 2) with a learning rate $\eta$. The following bound holds:
$\|W_{0}^{t}-W^{opt}\|\leq\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|W_{0}-W^{opt}\|.$
where $W_{0}$ is the initial model.
Proof 12
Let $lr$ denotes the total number of training iterations on $D_{0}$ over $t$ rounds of iteration. We define $w^{0}=W_{0}$, and after $lr$ rounds of local training on dataset $D_{0}$, the final model is denoted by $w^{lr}=W_{0}^{t}$. During the $i$-th round of training, the gradient descent update rule gives:
$w^{i+1}=w^{i}-\eta\nabla F(w^{i};D_{0}).$
Subtracting $W^{opt}$ from both sides, we obtain:
$w^{i+1}-W^{opt}=w^{i}-W^{opt}-\eta\nabla F(w^{i};D_{0}).$
Taking the norm and square on both sides, we have:
$\|w^{i+1}-W^{opt}\|^{2}$
$=$ $\|w^{i}-W^{opt}\|^{2}-2\eta\langle\nabla F(w^{i};D_{0}),w^{i}-W^{opt}\rangle$
$+\eta^{2}\|\nabla F(w^{i};D_{0})\|^{2}.$
By Assumption 1 and $\nabla F(W^{opt})=0$, we have:
$F(w^{i})+\langle\nabla F(w^{i};D_{0}),(W^{opt}-w^{i})\rangle$
$+\frac{\mu}{2}\|W^{opt}-w^{i}\|^{2}$ $\leq F(W^{opt})$
$F(W^{opt})+\frac{\mu}{2}\|w^{i}-W^{opt}\|^{2}$ $\leq F(w^{i})$
$\|\nabla F(w^{i};D_{0})-\nabla F(W^{opt})\|^{2}$ $\leq L^{2}\|w^{i}-W^{opt}\|^{2}$
Then, it satisfies the following property:
$-2\eta\langle\nabla F(w^{i};D_{0}),w^{i}-W^{opt}\rangle$ $\leq-2\eta\mu\|w^{i}-W^{opt}\|^{2}$
$\eta^{2}\|\nabla F(w^{i};D_{0})\|^{2}$ $\leq\eta^{2}L\|w^{i}-W^{opt}\|^{2}$
Then, we have
$\|w^{i+1}-W^{opt}\|^{2}\leq(1-2\eta\mu+\eta^{2}L^{2})\|w^{i}-W^{opt}\|^{2}$
$\|w^{i+1}-W^{opt}\|\leq\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\|w^{i}-W^{opt}\|$
Recursively applying this inequality over $lr$ rounds yields:
$\|w^{lr}-W^{opt}\|\leq\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|w^{0}-W^{opt}\|.$
Since $w^{0}=W_{0},w^{lr}=W_{0}^{t}$, the global model after $lr$ local training rounds satisfies:
$\|W_{0}^{t}-W^{opt}\|\leq\left(\sqrt{(1-2\eta\mu+\eta^{2}L^{2})}\right)^{lr}\|W_{0}-W^{opt}\|.$